### v3

# Rational motives on pro-algebraic stacks

Abstract
We define an $\infty$-category of rational motives for cofiltered limits of algebraic stacks, so-called pro-algebraic stacks. We show that it admits a $6$-functor formalism for certain classes of morphisms. Our main example is the stack of displays, closely related to the stack of Barsotti-Tate groups, seen as a pro-algebraic stack via its truncations. On the way we deal with base change along non-finite type morphisms such as non-finite Galois extensions.
1 Introduction
The study of algebraic cycles is important in algebraic geometry, number theory and representation theory. Over the years there have been many conjectures and results concerning algebraic cycles, that we could not list in its detail in this article. Not only have cycles on varieties been studied but also the study of equivariant cycles has become important, for example for the construction of higher Gross-Zagier formulas by Yun and Zhang (cf. [YZ17]).
For the study of equivariant cycles, Edidin and Graham constructed equivariant Chow groups (cf. [EG98]) following the idea of Totaro (cf. [Tot99]). In his PhD-thesis Brokemper gave a computation of the Chow ring of the stack of truncated Barsotti-Tate groups ([Bro18]). The main idea was to relate the Chow groups of the stack of truncated Barsotti-Tate groups to the stack of so-called truncated displays due to Lau ([Lau13]). Let us denote the stack of $n$-truncated displays by $\textup{Disp}_{n}$. This is an Artin stack over $\mathbb{F}_{p}$. For each $n\geq 1$, we get a map $\tau_{n}\colon\textup{Disp}_{n+1}\rightarrow\textup{Disp}_{n}$. Brokemper showed that this map admits a factorization
[…]
where $\alpha$ is an affine bundle and $\beta$ is a map of quotient stacks $\left[X/G\right]\rightarrow\left[X/H\right]$ induced by an extension of a unipotent group, i.e. an exact sequence
$0\rightarrow U\rightarrow G\rightarrow H\rightarrow 0,$
where $U$ is unipotent. By homotopy invariance, we therefore have $A^{*}(\textup{Disp}_{n},*)\cong A^{*}(\textup{Disp}_{1},*)$. One can show that $\textup{Disp}_{1}$ has a particular representation as a quotient stack, which allows Brokemper to compute its Chow groups explicitly. An interesting application of this result is the following. Let $\textup{BT}_{n}$ denote the stack of $n$-truncated Barsotti-Tate groups. Then there exists a smooth morphism $\phi_{n}\colon\textup{BT}_{n}\rightarrow\textup{Disp}_{n}$, that is compatible with truncations and is an equivalence on geometric points (cf. [Lau13]). This morphism induces an isomorphism on the level of rational Chow groups and an injection on the level of integral Chow groups (cf. [Bro18]). Rationally, this is clear as $\phi_{n}$ is a universal homeomorphism and thus induces an equivalence of motives. Thus, for any $n\in\mathbb{N}$, we have
$A^{*}(\textup{BT}_{n},*)_{\mathbb{Q}}\cong A^{*}(\textup{Disp}_{1},*)_{\mathbb{Q}}.$
Due to the work of Levine, we know that the Chow groups can be realized by motivic cohomology (cf. [Lev]). Rationally the same is true, even for equivariant Chow groups but let us be a bit more precise. For this paragraph, we assume that every Artin stack, and hence also every scheme, is of finite type over a field. The category of Beilinson motives satisfies étale descent (cf. [CD19]). Using this, we can follow the construction of Richarz and Scholbach to define $\textup{DM}(X)$, the $\infty$-category of rational motives on an Artin stack $X$ with smooth atlas $\tilde{X}$, via the formula
[…]
(cf. [RS20]). Equivalently, as shown in op. cit., we can define $\textup{DM}(X)$ via right Kan extension along the Yoneda embedding. This approach shows how to define $\textup{DM}(F)$ for any prestack $SSF\colon\textup{(Ring)}\rightarrow\SS$ (here $SS\SS$ denotes the $\infty$-category of Kan-complexes).
One particular question, that is the initial motivation for this article, is what happens with the motivic cohomology of the non-truncated stack of displays, i.e. the motivic cohomology of $\textup{Disp}\coloneqq\lim_{n}\textup{Disp}_{n}$. The motivic cohomology of Disp should be equal to the motivic cohomology of $\textup{Disp}_{1}$ since we enlarge $\textup{Disp}_{1}$ in a way that cannot be seen in $\mathbb{A}^{1}$-homotopy theory. However the computation of motivic cohomology via right Kan extension seems to more complicated. Instead of this approach, we will give an alternative definition of DM for cofiltered limits of algebraic stacks of finite type, so-called pro-algebraic stacks.
So let us fix a filtered category ${\mathcal{I}}$ and for the sake of this introduction let us fix a field $k$ and $S\coloneqq\operatorname{Spec}(k)$.
Definition 1.
A $-{\mathcal{I}}$-algebraic stack is a map $x\colon{\mathcal{I}}^{\operatorname{op}}\rightarrow\textup{PreStk}_{S}$, i.e. a cofiltered systems of prestacks, such that each $X_{i}\coloneqq x(i)$ is an Artin stack locally of finite type over $S$. We denote such stacks as $(x,X)$, where $X\coloneqq\lim_{{\mathcal{I}}}x$. The transition maps are denoted by $x_{ij}\colon X_{j}\rightarrow X_{i}$ for any map $i\rightarrow j\in{\mathcal{I}}$ (this is an abusive notation but for convenience the ambiguity can be neglected). We say that $(x,X)$ is
  1. 1.
    classical if $x_{ij}$ is affine for all $i\rightarrow j\in{\mathcal{I}}$,
  2. 2.
    monoidal if $x_{ij}$ is smooth for all $i\rightarrow j\in{\mathcal{I}}$, and
  3. 3.
    strict if $x_{ij}^{*}\colon\textup{DM}(X_{i})\rightarrow\textup{DM}(X_{j})$ is fully faithful for all $i\rightarrow j\in{\mathcal{I}}$.
A morphism $f\colon(x,X)\rightarrow(y,Y)$ of $-{\mathcal{I}}$-algebraic stacks is a morphism of diagrams. We denote by $f_{i}\colon X_{i}\rightarrow Y_{i}$ the induced map. We call a morphism $f\colon(x,X)\rightarrow(y,Y)$ of $-{\mathcal{I}}$-algebraic stacks cartesian if each square
[…]
is a pullback diagram. We call a morphism $f\colon(x,X)\rightarrow(y,Y)$ of $-{\mathcal{I}}$-algebraic stacks adjointable if the morphism $y_{ij}^{*}f_{i*}\rightarrow f_{j*}x_{ij}^{*}$ induced by each square above is an equivalence.
Let us mention $2$ examples of pro-algebraic stacks.
Example 2.
  1. (1)
    Let $k$ be a field and $L/k$ be a (not necessarily finite) Galois extension. Then the system of finite Galois extensions ${\mathcal{G}}_{k}^{L}$ of $k$ containing $L$ induces a classical monoidal pro-${\mathcal{G}}_{k}^{L}$-algebraic stack. Similarly, via base change, we can define for any algebraic $k$-stack $f\colon{\mathfrak{X}}\rightarrow\operatorname{Spec}(k)$ a pro-${\mathcal{G}}_{k}^{L}$-algebraic stack $(q_{{\mathfrak{X}}},{\mathfrak{X}}_{L})$ that comes with a cartesian map to $(q,\operatorname{Spec}(L))$.
  2. (2)
    The $-\mathbb{N}$-algebraic stack $(\tau,\textup{Disp})$ is by the discussion above strict and monoidal but not classical.
Example (1) above will show that our formalism of motives on pro-algebraic stacks allows us to work with non-finite type morphisms of schemes and get for example a smooth base change in this setting. Example (2) was discussed in the beginning of this introduction and will show how to define motives for stacks that are not necessarily algebraic.
Let us also give $2$ interesting examples of adjointable morphisms.
Example 3.
  1. (1)
    Let $f\colon(x,X)\rightarrow(y,Y)$ be a cartesian morphism of monoidal $-{\mathcal{I}}$-algebraic stacks, then by smooth base change $f$ is adjointable.
  2. (2)
    Let $f\colon(x,X)\rightarrow(y,Y)$ be a morphism of $-{\mathcal{I}}$-algebraic stacks. Assume that each $y_{ij*}$ is fully faithful and $(x,X)$ is strict, then simple computation shows that $f$ is adjointable. An example of such a situation is the structure map $(\tau,\textup{Disp})\rightarrow\operatorname{Spec}(\mathbb{F}_{p})$, where $(\tau,\textup{Disp})$ is as in Example 2.
Definition 4.
We define the $\infty$-category of rational motives on a $-{\mathcal{I}}$-algebraic stack $(x,X)$ to be
$\textup{DM}(x,X)\coloneqq\mathop{\textup{colim}}_{i\in{\mathcal{I}},x^{*}}\textup{DM}(X_{i}).$
Note that if $(x,X)$ is classical, then $X$ is in fact representable by an algebraic stack and we have $\textup{DM}(x,X)\simeq\textup{DM}(X)$ (see Proposition 3.1). Also, with this definition, it is clear that $\textup{DM}(x,X)$ admits a closed symmetric monoidal structure and is compactly generated. One of the main motivations behind our choice of definition is that we obtain a (motivic) $6$-functor formalism for certain classes of morphisms.
Theorem 5.
Let $f\colon(x,X)\rightarrow(y,Y)$ be a morphism of monoidal $-{\mathcal{I}}$-algebraic stacks. Then there is an adjunction
${f^{*}\colon\textup{DM}(y,Y)}$${\textup{DM}(x,X)\colon f_{*}}$
functorial in $(x,X)$ and $(y,Y)$, where $f^{*}$ is symmetric monoidal and preserves compact objects. Moreover, the functor DM together with $f^{*}$ defines a pullback formalism, i.e. if $f$ is smooth then $f^{*}$ admits a left adjoint $f_{\sharp}$ that satisfies base change and projection formula.
Assume now that ${\mathcal{I}}\simeq\mathbb{N}_{0}$ and $f$ is adjointable or that ${\mathcal{I}}$ is arbitrary and $f$ is cartesian. Then we can upgrade $f^{*}$ to be part of a $6$-functor formalism in the following sense. There exists a functorial adjunction
${f_{!}\colon\textup{DM}(x,X)}$${\textup{DM}(y,Y)\colon f^{!}.}$
Further, the adjunctions above satisfy the following properties.
  1. (F1)
    If $f$ is proper, we have a canonical equivalence $f_{*}\simeq f_{!}$.
  2. (F2)
    The functors $f_{!}$ and $f^{!}$ commute with arbitrary colimits.
  3. (F3)
    (Base change) For any pullback square
    ${(w,W)}$${(z,Z)}$${(x,X)}$${(y,Y)}$$\scriptstyle{f^{\prime}}$$\scriptstyle{g^{\prime}}$$\scriptstyle{g}$$\scriptstyle{f}$
    of monoidal pro-algebraic stacks with morphisms satisfying (i) or (ii), the exchange morphisms
    $f^{*}g_{!}\rightarrow g^{\prime}_{!}f^{\prime*}$
    $g^{!}f_{*}\rightarrow f^{\prime}_{*}g^{\prime!}$
    are equivalences.
  4. (F4)
    (Projection formula) The exchange transformation $f_{!}(-\otimes f^{*}(-))\rightarrow f_{!}(-)\otimes-$ is an equivalence.
  5. (M1)
    (Homotopy invariance) Let us define $(q,\mathbb{A}^{1}_{X})$ via the system $n\mapsto\mathbb{A}^{1}_{X_{i}}$. Then $p^{*}$ is fully faithful, where $p\colon(q,\mathbb{A}^{1}_{X})\rightarrow(x,X)$ denotes the projection.
  6. (M2)
    (T-stability) Let $V_{0}\rightarrow X_{0}$ be a vector bundle and let $p\colon(v,V)\rightarrow(x,X)$ be the induced morphism of pro-algebraic stacks. The zero section $s_{0}\colon X_{0}\rightarrow V_{0}$ induces a section $s\colon(x,X)\rightarrow(v,V)$. Then the Thom-map $\textup{Th}(p,s)\coloneqq\textup{Th}(V_{0})\coloneqq p_{\sharp}s_{*}$ is an equivalence of DG-categories.
  7. (M3)
    (Purity) We have a canonical equivalence $f^{!}\simeq\operatorname{\textup{Th}}(\Omega_{f_{0}})1_{X}\otimes f^{*}$.
  8. (M4)
    (Orientation) Let $(q,\operatorname{\mathbb{G}_{m}}_{,X})$ be the pro-algebraic stack defined by $n\mapsto\operatorname{\mathbb{G}_{m}}_{,X_{i}}$. We naturally obtain a projection $p\colon(q,\operatorname{\mathbb{G}_{m}}_{,X})\rightarrow(x,X)$. For $M\in\textup{DM}(X)$, we set $M(1)\coloneqq\operatorname{\textup{cofib}}(p_{\sharp}p^{*}M\rightarrow M)[-1]$. Let $V_{0}\rightarrow X_{0}$ be a vector bundle of rank $n$. Then there is a functorial equivalence $\textup{Th}(V_{0})1_{X}\simeq 1_{X}(n)[2n]$.
  9. (M5)
    (Localization) Let $i\colon(z,Z)\hookrightarrow(x,X)$ be cartesian closed immersion of pro-algebraic stacks with open complement $j\colon(u,U)\rightarrow(x,X)$. Then there exist fiber sequences
    $i_{!}i^{!}\rightarrow\operatorname{id}\rightarrow j_{*}j^{*}$
    $j_{!}j^{!}\rightarrow\operatorname{id}\rightarrow i_{*}i^{*}.$
Even though the proof is formal, this theorem allows us to explicitly compute the motive of (strict) $-{\mathcal{I}}$-algebraic stacks and their motivic cohomology. For example, if $(x,X)$ is a strict monoidal $-{\mathcal{I}}$-algebraic stack with a map $f\colon(x,X)\rightarrow Y$, where $Y$ is an algebraic stack seen as a constant diagram, then
$f_{!}f^{!}1_{Y}\simeq f_{0!}f^{!}_{0}1_{Y},$
as expected if ${\mathcal{I}}$ has an initial object 00.
Remark 6.
Let us quickly mention that we develop the theory of rational motives not only for pro-algebraic stacks. Theorem 5 works similarly for any (motivic) six functor formalisms that are compactly generated if we restrict us to $-{\mathcal{I}}$-schemes (analogously defined to the stack case). If further the (motivic) six functor formalism satisfies étale descent, then we this also holds for $-{\mathcal{I}}$-algebraic stacks.
Let us define absolute motivic cohomology as follows.
Definition 7.
Let $(x,X)$ be a monoidal $-{\mathcal{I}}$-algebraic stack and $n,m\in\mathbb{Z}$. Then we define the absolute motivic cohomology of $(x,X)$ in degree $(n,m)$ by
$H^{n,m}(X,\mathbb{Q})\coloneqq\operatorname{Hom}_{\textup{DM}(x,X)}(1_{(x,X)},1_{(x,X)}(n)[m]),$
where we define Tate-twists according to Theorem 5.
We will see in Proposition 3.10 that if $(x,X)$ is classical and $X$ is represented by an algebraic stack locally of finite type over $S$, then $H^{n,m}(X,\mathbb{Q})\cong\operatorname{Hom}_{\textup{DM}(X)}(1_{X},1_{X}(n)[m])$ recovering the usual definition. Even in the non-classical case, following Brokemper’s thesis (cf. [Bro16]), we will see in Example 3.12 that if $k$ is a field, $L$ and algebraic extension and $Z$ a $k$-scheme, then $H^{n,m}(Z_{L},\mathbb{Q})\cong A^{n}(Z_{L},2n-m)_{\mathbb{Q}}$ as expected.
Let us now come back to the setting of the beginning of this introduction, i.e. Example 2 (2). Recall, that the truncation $\tau_{n}\colon\textup{Disp}_{n+1}\rightarrow\textup{Disp}_{n}$ is a composition of an affine bundle and a map of quotient stacks induced by an extension of a unipotent group. Then the following observation follows immediately from construction and the author’s computations of Tate motives on quotients by Frobenius conjugation (cf. [Yay23a])
Theorem 8 (4.10).
Let $f\colon\textup{Disp}\rightarrow\operatorname{Spec}(\mathbb{F}_{p})$ denote the structure map. Then $f_{*}f^{*}1_{\mathbb{F}_{p}}\simeq f_{1*}f_{1}^{*}1_{\mathbb{F}_{p}}$ and $f_{\sharp}f^{*}1_{\mathbb{F}_{p}}\simeq f_{1\sharp}f_{1}^{*}1_{\mathbb{F}_{p}}$ are contained in the ind-completion of the full stable subcategory of $\textup{DM}(S)$ generated by Tate-motives.
The equivalence $f_{*}f^{*}1_{\mathbb{F}_{p}}\simeq f_{1*}f_{1}^{*}1_{\mathbb{F}_{p}}$ stated in the above theorem also allows us to compute the motivic cohomology of Disp after truncation (this has been done by Brokemper, cf. [Bro18]).
Structure of this article
We start by defining pro-algebraic stacks and listing some examples. We define the $\infty$-category of rational motives on pro-algebraic stacks.
Afterward, we prove our main theorem, that on certain classes of morphisms, this definition yields a (motivic) $6$-functor formalism. We do this in a very general setting considering pullback formalisms, coefficient systems and abstract $6$-functor formalisms.
Lastly, we apply our theory to the three examples mentioned in the introduction.
Acknowledgement
This project started after a talk with Torsten Wedhorn about motives on limits of algebraic stacks and I would like to thank him for various discussions and his interest in this project. Further, I would like to thank Rizacan Ciloglu, David G. Davis, Martin Gallauer, Benjamin Hennion, Ryomei Iwasa, Hyungseop Kim, Eike Lau, Timo Richarz, Jakob Scholbach and Thibaud van den Hove for helpful discussions.
This project was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - project number 524431573, the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) TRR 326 Geometry and Arithmetic of Uniformized Structures, project number 444845124 and by the LOEWE grant ‘Uniformized Structures in Algebra and Geometry’.
1.1 Assumptions and notations
In the following, we want to fix some categorical and algebraic notation that is used throughout this article.
Categorical notation
Throughout, we fix some inaccessible regular cardinal $\kappa$. By small, we will mean $\kappa$-small. We will freely work with $(\infty,1)$-categories in the sense of [Lur09] and if we write $\infty$-category, we will always mean $(\infty,1)$-category. When we say category, we always mean $1$-category and view it as an $\infty$-category via the nerve functor.
  1. $\bullet$
    Let $K$ be a poset. Then we view $K$ as an $\infty$-category with morphisms induced by the ordering.
  2. $\bullet$
    If $F\colon C\rightarrow D$ is a functor of $\infty$-categories that admits a right adjoint $G$, then we will denote the adjunction of $F$ and $G$ by the symbol $F\dashv G$.
  3. $\bullet$
    The $\infty$-category of small $\infty$-groupoids is denoted by $SS\SS$.
  4. $\bullet$
    The $\infty$-category of small $\infty$-categories is denotes by $\textup{Cat}_{\infty}$.
  5. $\bullet$
    The $\infty$-category of presentable stable $\infty$-categories with continuous functors, i.e. filtered colimit preserving functors, is denoted by $\textup{Pr}^{\textup{L}}_{\textup{st}}$.
  6. $\bullet$
    Let ${\mathcal{D}}(\mathbb{Q})$ denote the category derived category of $\mathbb{Q}$-vector spaces. This is an algebra object in $\textup{Pr}^{\textup{L}}_{\textup{st}}$. The $\infty$-category $\textup{DGCat}_{\textup{cont}}$ denotes the $\infty$-category of ${\mathcal{D}}(\mathbb{Q})$-modules in $\textup{Pr}^{\textup{L}}_{\textup{st}}$.
  7. $\bullet$
    Throughout this article, we will deal with diagrams $X\in\operatorname{Fun}(K,{\mathcal{C}})$, where $K$ and ${\mathcal{C}}$ are $\infty$-categories. For any $k\in K$, we will denote by $X_{k}$ the value of $X$ at $k$ and for any morphism $\alpha\colon k\rightarrow l$ in $K$, we will denote the induced morphism in ${\mathcal{C}}$ by $X_{\alpha}$. We will often write $X_{kl}$ whenever the morphism $k\rightarrow l$ is clear from the context. This abusive notation is for convenience of the reader.
    If $f\colon X\rightarrow Y$ is a morphisms of diagrams in $\operatorname{Fun}(K,{\mathcal{C}})$, then we write $f_{k}\colon X_{k}\rightarrow Y_{k}$ for the morphism in ${\mathcal{C}}$ associated to $k\in K$.
Algebraic notation
Throughout, let $B$ be an excellent noetherian scheme of dimension $\leq 2$ and let $S$ be a regular scheme of finite type over $B$.
  1. $\bullet$
    An Artin stack is an algebraic stack in the sense of [Sta22].
  2. $\bullet$
    Let $X$ be an $S$-scheme of finite type. By $\textup{DM}(X)$, we denote the DG-category of Beilinson-motives (cf. [CD19, §14]). Roughly, these are $\mathbb{A}^{1}$-local étale sheaves on $X$ with values in ${\mathcal{D}}(\mathbb{Q})$ such that $\mathbb{P}^{1}$ is $\otimes$-invertible.
2 (Motivic) six functor formalisms on pro-algebraic stacks
In this article, we want to extend the 6-functor formalism to limits of Artin stacks, for example the stack of Barsotti-Tate groups. The stack of such groups can be obtained as a limit of its truncations. We will explain later that these truncation maps are given by affine bundles (see also [Bro18]). In particular, if we have a 6-functor formalism for motives on such stacks, we can analyze their motive by truncating. All of this will be made more precise later on. But first let us establish a general theory on inverse systems of Artin stacks that may have non-representable transition maps.
Before we begin with our main constructions and definitions, we want to recall the theory of rational motives on Artin stacks. We will follow the construction of Richarz-Scholbach to obtain the DG-category of rational motives on prestacks (cf. [RS20, Def. 2.2.1]). Later on we will specialize this to our case. In the following, we will omit some details and refer to loc. cit. for an explicit construction.
We denote by $\textup{AffSch}^{\textup{ft}}_{S}$ the category of affine schemes over $S$. As $\textup{AffSch}^{\textup{ft}}_{S}$ is essentially small we replace $\textup{AffSch}^{\textup{ft}}_{S}$ by a small skeleton of objects of interest in this article. We denote by $\textup{AffSch}^{\kappa}_{S}$ the pro-$\kappa$-completion of $\textup{AffSch}^{\textup{ft}}_{S}$. Again, this category is small (cf. [KS06, Lem. 6.1.2]).
A prestack over $S$ is a functor from $\textup{AffSch}^{\kappa}_{S}$ to $\infty$-groupoids and we denote the $\infty$-category of prestacks over $S$ by $\textup{PreStk}_{S}$.
The functor $\textup{DM}\colon X\mapsto\textup{DM}(X)$ denotes the functor of Beilinson motives on finite type $S$-schemes. We extend DM to prestacks in the following way. First we left Kan extend DM from $\textup{AffSch}^{\textup{ft}}_{S}$ to $\textup{AffSch}^{\kappa}_{S}$ and then right Kan extend along the Yoneda embedding $\textup{AffSch}^{\kappa}_{S}\hookrightarrow\textup{PreStk}_{S}$, where the transition maps are given by *-pullbacks. This construction yields a functor
$\textup{DM}^{*}\colon\textup{PreStk}_{S}\rightarrow\textup{DGCat}_{\textup{cont}}.$
If instead of *-pullback, we use the !!!-pullback, we denote this functor by $\textup{DM}^{!}$. However, the restriction to schemes locally of finite type over $S$ yields equivalent descriptions of $S$. This is due to the fact that in our setting DM for schemes satisfies h-descent (cf. [CD19]). Thus, both $\textup{DM}^{!}$ and $\textup{DM}^{*}$ can be computed via descent. This follows as locally finite type schemes admit an open covering by affines of finite type and !!!- resp. *-pullback agree up to shift and twist. Note that the same argumentation can be used to see that $\textup{DM}^{*}(X)\simeq\textup{DM}^{!}(X)$ for any algebraic space, as the diagonal of an algebraic space is representable by a scheme.
Using étale descent, we can glue the 6-functor formalism of DM to the restriction of $\textup{DM}^{*}$ to algebraic spaces, as in [Kha19] or equivalently by using the DESCENT algorithm of Liu-Zheng (cf. [LZ17]). Similar constructions can also be found in [RS20], where ind-Artin stacks and ind-schemes are considered.
Finally, we can now apply the same arguments to glue the 6-functor formalism to the restriction of $\textup{DM}^{*}$ to Artin stacks locally of finite type. As before the existence of a smooth cover on an Artin stack shows $\textup{DM}^{!}(X)\simeq\textup{DM}^{*}(X)$ for any Artin stack $X$ locally of finite type over $S$.
To ease notation, we will from here on denote by DM the extension $\textup{DM}^{*}$. For the !!!-version, we will keep the notation $\textup{DM}^{!}$.
Notation 2.1.
Throughout this article, we will denote by ${\mathcal{I}}$ a filtered $\infty$-category.
Definition 2.2.
Let $x\colon{\mathcal{I}}^{\operatorname{op}}\rightarrow\textup{PreStk}_{S}$ be a filtered system of prestacks and let $X\coloneqq\lim x$ be the associated limit prestack. We call the tuple $(x,X)$ a $-{\mathcal{I}}$-algebraic stack, if $x(i)$ is an algebraic stack locally of finite type over $S$ for all $i\in{\mathcal{I}}$.
If $(x,X)$ is a pro-algebraic stack, then we denote by $X_{i}\coloneqq x(i)$ and the transition maps $X_{j}\rightarrow X_{i}$ by $x_{ij}$, for $i\rightarrow j\in{\mathcal{I}}$.
A morphism $f\colon(x,X)\rightarrow(y,Y)$ of pro-algebraic stacks is a morphism $f_{\bullet}\colon x\rightarrow y$. If ${\bf P}$ is a property of a morphism of algebraic stacks, we will say that $f$ has property ${\bf P}$ if each $f_{i}\colon X_{i}\rightarrow Y_{i}$ has property ${\bf P}$.
We denote by $\textup{ProAlg}^{{\mathcal{I}}}$ the $\infty$-category of $-{\mathcal{I}}$-algebraic stacks.
If $X$ is an algebraic stack, we will abuse notation to denote the constant diagram $n\mapsto X$ (here the transition maps are given by $\operatorname{id}_{X}$) with $X$ itself.
Definition 2.3.
If $f\colon(x,X)\rightarrow(y,Y)$ is a morphism of $-{\mathcal{I}}$-algebraic stacks, then we call $f$ cartesian if each square
[…]
is a pullback diagram. We call a morphism $f\colon(x,X)\rightarrow(y,Y)$ of $-{\mathcal{I}}$-algebraic stacks adjointable if the morphism $y_{ij}^{*}f_{i*}\rightarrow f_{j*}x_{ij}^{*}$ induced by each square above is an equivalence.
Example 2.4.
Let us give two important examples of adjointable morphisms of $-{\mathcal{I}}$-algebraic stacks.
  1. (1)
    Let $f\colon(x,X)\rightarrow(y,Y)$ be a cartesian morphism of monoidal $-{\mathcal{I}}$-algebraic stacks, then by smooth base change $f$ is adjointable.
  2. (2)
    Let $f\colon(x,X)\rightarrow(y,Y)$ be a morphism of $-{\mathcal{I}}$-algebraic stacks. Assume that each $y_{ij*}$ is fully faithful (e.g. when $Y$ is equivalent to a constant diagram of induced by an algebraic stack) and $(x,X)$ is strict, then simple computation shows that $f$ is adjointable.
Definition 2.5.
Let $(x,X)$ be a $-{\mathcal{I}}$-algebraic stack. Then we say that $(x,X)$ is
  1. 1.
    classical if $x_{ij}$ is affine for all $i\rightarrow j\in{\mathcal{I}}$,
  2. 2.
    monoidal if $x_{ij}$ is smooth for all $i\rightarrow j\in{\mathcal{I}}$, and
  3. 3.
    strict if $x_{ij}^{*}\colon\textup{DM}(X_{i})\rightarrow\textup{DM}(X_{j})$ is fully faithful for all $i\rightarrow j\in{\mathcal{I}}$.
Example 2.6.
Let $(x,X)$ be a $-{\mathcal{I}}$-algebraic stack and assume that each $x_{ij}$ is an affine bundle for all $i\rightarrow j\in{\mathcal{I}}$, i.e. Zariski locally isomorphic to an affine space, then $(x,X)$ is a classical monoidal strict $-{\mathcal{I}}$-algebraic stack.
Example 2.7.
[…]
The above can be generalized to the following case. We can replace $\operatorname{Spec}(k)$ by $S$ and assume that $S$ is connected. Let $\bar{s}$ be a geometric point of $S$. Then we can define $F_{S}$ as the category of pointed étale covers over $(S,\bar{s})$. The construction above can be repeated in this case for any smooth Artin stack ${\mathfrak{X}}$ over $S$.
Now let us define the DG-category of rational motives associated to a pro-algebraic stack.
Definition 2.8.
Let $(x,X)$ be a $-{\mathcal{I}}$-algebraic stack. Then we define the $\infty$-category of rational motives on $(x,X)$ as
$\textup{DM}(x,X)\coloneqq\mathop{\textup{colim}}_{i\in{\mathcal{I}},x^{*}}\textup{DM}^{*}(X_{i}),$
where the colimit is taken in $\textup{DGCat}_{\textup{cont}}$.
One might wonder why we chose $\textup{DM}^{*}$ and not $\textup{DM}^{!}$. Our main reason is convenience. For monoidal $-{\mathcal{I}}$-algebraic stacks, we conjecture that we have an equivalence between both possible definitions. But by working with $\textup{DM}^{*}$, monoidality of $\textup{DM}(x,X)$ follows straightforwardly. We want to remark that with this definition of $\textup{DM}(x,X)$, we recover the usual $\textup{DM}(X)$ if $(x,X)$ is classical monoidal.
In rest of this section is we want to prove Theorem 5.
2.1 Pullback formalisms
We will first define pullback formalisms on schemes after Drew-Gallauer (cf. [DG22]). Afterward, we want to show that DM on $-{\mathcal{I}}$-algebraic stacks defines a pullback formalism. We will analyze such structures more generally in this subsection without using $-{\mathcal{I}}$-algebraic stacks.
Definition 2.9 ([DG22]).
Let ${\mathcal{E}}$ be an $\infty$-category that admits finite limits and let us fix a class of morphisms $P\subseteq\operatorname{Fun}(\Delta^{1},{\mathcal{E}})$ that is closed under pullback and equivalences and composition. A pullback formalism on ${\mathcal{E}}$ with respect to $P$ is a functor
${\mathcal{M}}^{*,\otimes}\colon{\mathcal{E}}^{\operatorname{op}}\rightarrow\textup{CAlg}(\textup{Pr}^{L,\otimes}),\quad f\mapsto f^{*}\quad\textup{(we denote the right adjoint of $f^{*}$ by $f_{*}$)}$
satisfying the following conditions.
  1. (1)
    If a morphism $f\colon X\rightarrow Y$ of $S$ lies in $P$, then there exists a left adjoint $f_{\sharp}$ of $f^{*}$.
  2. (2)
    For each pullback square
    ${X\times_{Z}Y}$${X}$${Y}$${Z}$$\scriptstyle{g}$$\scriptstyle{f}$$\scriptstyle{f}$$\scriptstyle{g}$
    in ${\mathcal{S}}$ such that $f$ is in $P$, the exchange transformation
    $f^{\prime}_{\sharp}g^{\prime*}\rightarrow f^{\prime}_{\sharp}g^{\prime*}f^{*}f_{\sharp}\rightarrow f^{\prime}_{\sharp}f^{\prime*}g^{*}f_{\sharp}\rightarrow g^{*}f_{\sharp}$
    is equivalence.
  3. (3)
    For each morphism $f\colon X\rightarrow Y$ in $P$ the exchange transformation
    $f_{\sharp}(M\otimes_{{\mathcal{M}}^{*,\otimes}(X)}f^{*}N)\rightarrow f_{\sharp}M\otimes_{{\mathcal{M}}^{*,\otimes}(Y)}N$
    is an equivalence, for any $M\in{\mathcal{M}}^{*,\otimes}(X)$ and $N\in{\mathcal{M}}^{*,\otimes}(Y)$.
Let us remark that this definition of a pullback formalism differs slightly from the definition of op. cit.. This adjustment is necessary for our means and a for a more general overview we refer to op. cit..
Remark 2.10.
For property (2) in Definition 2.9 we may equivalently ask for the *-exchange morphism $f^{*}g_{*}\rightarrow g^{\prime}_{*}f^{\prime*}$ to be an equivalence (cf. [Ayo07, Prop. 1.1.9]).
Example 2.11.
Throughout this article, we will be mainly interested in the case where ${\mathcal{E}}\simeq\textup{Stk}^{\textup{lft}}_{S}$, where $\textup{Stk}^{\textup{lft}}_{S}$ denotes the category of locally of finite type Artin stacks over $S$ (recall $S$ from our notations), $P$ the class of smooth morphisms and ${\mathcal{M}}^{*,\otimes}=\textup{DM}_{\mathbb{Q}}$ with the *-pullback. If ${\mathcal{E}}$ is the category of finite type $S$-schemes, instead of $\textup{DM}_{\mathbb{Q}}$, we could also work $\textup{DM}_{\mathbb{Z}}$ or even with the stable homotopy category SH.
Definition 2.12.
A pullback formalism ${\mathcal{M}}^{*,\otimes}$ is called compactly generated if ${\mathcal{M}}^{*,\otimes}$ factors through the functor $\textup{CAlg}(\textup{Pr}^{L,\otimes}_{\omega})$, where $\omega$ denotes the cardinality of the natural numbers.
Example 2.13.
Recall that in the assumptions 1.1, we fixed a regular scheme $S$ of finite type over a excellent noetherian scheme $B$ of dimension $\leq 2$. By [CD19, Thm. 16.1.4, Cor. 6.2.2] the $\infty$-category $\textup{DM}(X,\mathbb{Q})$ is compactly generated for any finite type $S$-scheme and $f^{*}$ preserves compact objects. In particular, following Example 2.11, $\textup{DM}_{\mathbb{Q}}$ induces a compactly generated pullback formalism.
Notation 2.14.
For this subsection, we fix an $\infty$-category ${\mathcal{E}}$ that admits finite limits, a class of morphisms $P_{{\mathcal{E}}}$ that is stable under pullbacks, compositions and equivalences. Further, we fix a compactly generated pullback formalism ${\mathcal{M}}^{*,\otimes}$ on ${\mathcal{E}}$ with respect to $P_{{\mathcal{E}}}$ and simply denote it by ${\mathcal{M}}^{\otimes}$.
Example 2.15.
Let us give $2$ examples for ${\mathcal{E}}$, $P_{{\mathcal{E}}}$ and ${\mathcal{M}}^{\otimes}$ that satisfy our assumptions. Recall $S$ from our assumptions and notations 1.1.
  1. (1)
    Let ${\mathcal{E}}=\textup{Sch}_{S}^{\textup{lft}}$ be the category of $S$-schemes locally of finite type and $P_{{\mathcal{C}}}$ the class of smooth morphisms in $\textup{Sch}_{S}$. Then $\textup{DM}_{\mathbb{Z}}$ together with *-pullbacks satisfies our assumptions.
  2. (2)
    Let us consider $\textup{DM}_{\mathbb{Q}}$ together with *-pullbacks. Then $\textup{DM}_{\mathbb{Q}}$ satisfies étale descent and standard arguments show that we can take ${\mathcal{E}}=\textup{Stk}^{\textup{lft}}_{S}$ and $P_{{\mathcal{E}}}$ to be the class of smooth morphisms (cf. [LZ12]).
We will start by extending ${\mathcal{M}}^{\otimes}$ to diagrams in $K$, where $K$ is a simplicial set, via taking colimits. By the remark below $\textup{CAlg}(\textup{Pr}^{L,\otimes}_{\omega})$ admits colimits.
Remark 2.16.
Let us remark that $\Pr^{L}_{\omega}$ is presentable by [Lur17, Lem. 5.3.2.9], thus so is $\textup{Mod}_{\operatorname{Sp}}(\textup{Pr}^{L,\otimes}_{\omega})$ by [Lur17, Cor. 4.2.3.7] and hence also $\textup{CAlg}(\textup{Pr}^{L,\otimes}_{\omega})$ by [Lur17, Prop. 3.2.3.5].
Let $K$ be a simplicial set. Then ${\mathcal{M}}$ induces a functor
$\operatorname{Fun}(K,{\mathcal{E}}^{\operatorname{op}})\rightarrow\operatorname{Fun}(K,\textup{CAlg}(\textup{Pr}^{L,\otimes}_{\omega})).$
Composing this functor with the colimit functor, we obtain
${\mathcal{M}}^{\otimes}_{K}\colon\operatorname{Fun}(K,{\mathcal{E}}^{\operatorname{op}})\rightarrow\textup{CAlg}(\textup{Pr}^{L,\otimes}_{\omega}).$
If $(x,X)$ is a $-{\mathcal{I}}$-algebraic stack, we see that the following notation agrees Definition 2.8.
Notation 2.17.
Let ${\mathcal{N}}$ be a pullback formalism on locally of finite type Artin $S$-stacks with respect to smooth morphisms. Let $(x,X)$ be a $-{\mathcal{I}}$-algebraic stack. Then we define
${\mathcal{N}}(x,X)\coloneqq{\mathcal{N}}_{{\mathcal{I}}}^{\otimes}(x).$
Proposition 2.18.
Let $K$ be a filtered $\infty$-category. Let ${\mathcal{P}}_{K}$ denote the class of morphisms in $\operatorname{Fun}(K,{\mathcal{E}})$ such that $f\in\operatorname{Fun}(K,{\mathcal{E}})$ belongs to ${\mathcal{P}}_{K}$ if and only if each $f_{k}$ belongs to $P_{{\mathcal{E}}}$ and each square in ${\mathcal{M}}_{K}(f)$ is right adjointable, i.e. for any morphism $i\rightarrow j\in K$ the induced morphism $f_{i}^{*}x_{ij*}\rightarrow y_{ij*}f_{j}^{*}$ is an equivalence (cf. [Lur17, Def. 4.7.4.13]).
Then ${\mathcal{M}}_{K}$ is a pullback formalism on $\operatorname{Fun}(K,{\mathcal{E}})$ for the class ${\mathcal{P}}_{K}$.
Proposition 2.18 is one of the crucial points in extending the motivic six functor formalism on DM for Artin-stacks to $-{\mathcal{I}}$-algebraic stacks. But before we can prove this proposition, we need some additional lemmas.
Lemma 2.19.
Let $K$ a filtered $\infty$-category. Let $F\colon{\mathcal{E}}\rightarrow\Pr^{L}_{\omega}$ be a functor. Denote by $F_{K}$ the composition of $\operatorname{Fun}(K,F)$ with the colimit functor. Let $X$ be in $\operatorname{Fun}(E,{\mathcal{C}})$. Then we have
$F_{K}(X)\simeq\textup{Ind}(\mathop{\textup{colim}}_{i\in K,x^{*}}F(X_{i})^{c})$
and any $M\in F_{K}(X)$ is compact if and only if there exists an $i\in K$ and $M_{i}\in F(X_{i})^{c}$ that lifts $M$.
Proof.
The first assertion follows from (cf. [GR17, Cor. 7.2.7]), where the colimit on the right is taken in $\textup{Cat}_{\infty}$.
Note that the filtered colimit of idempotent complete $\infty$-categories is idempotent complete (cf. [Lur09, Cor. 4.4.5.21]). Thus, any compact object comes from an object of the colimit on the right hand side (cf. [Lur09, Lem. 5.4.2.4] and [Lur21, 02LG]). ∎
Lemma 2.20.
Let $K$ be a $\infty$-category. Let $F\colon{\mathcal{E}}\rightarrow\Pr^{L}_{\omega}$ be a functor. Denote by $F_{K}$ the composition of $\operatorname{Fun}(K,F)$ with the colimit functor. Then we have the following.
  1. (1)
    Let $X\in\operatorname{Fun}(K,{\mathcal{E}})$. We denote by $X^{\operatorname{op}}$ the opposite functor. Further, let us denote by $F_{*}\colon{\mathcal{E}}^{\operatorname{op}}\rightarrow\Pr^{R}$ the functor induced by the equivalence $\Pr^{L}\simeq(\Pr^{R})^{\operatorname{op}}$ (cf. [Lur09, Cor. 5.5.3.4]). Then we have
    $F_{K}(X)\simeq\lim_{\Pr^{R}}F_{*}\circ X^{\operatorname{op}},$
    where the equivalence is in $\textup{Cat}_{\infty}$. More precisely, $F_{K}$ is equivalent to the right Kan extension of $F_{*}$ along $\delta^{\operatorname{op}}$.
  2. (2)
    If $K$ be a filtered $\infty$-category, then for any $k_{0},k_{1}\in K$ the composition
    $F(X_{k_{0}})\xrightarrow{\textup{ins}_{k_{0}}^{X}}F_{K}(X)\simeq\lim_{\Pr^{R}}F_{*}\circ X\xrightarrow{p_{k_{1}}^{X}}{\mathcal{M}}(X_{k_{1}})$
    is canonically equivalent to
    $\mathop{\textup{colim}}_{k\in K,\alpha\colon k_{0}\rightarrow k,\beta\colon k_{1}\rightarrow k}F_{*}(X_{\beta})\circ F_{K}(X_{\alpha})$
    where the colimit is taken in $\operatorname{Fun}_{\Pr^{L}}(F(X_{k_{0}}),F(X_{k_{1}}))$.
Proof.
The first assertion follows from [Lur09, Cor. 5.5.3.4] (see also Remark 2.21 below). A proof for (2) is given by Gaitsgory in the setting of DG-categories (cf. [Gai12, Lem. 1.3.6] - note that the proof is completely analogous in this slightly more general setting). ∎
Remark 2.21.
The equivalence in Lemma 2.20 for $X\in\operatorname{Fun}(K,{\mathcal{E}})$ is induced via the following process (cf. [Gai12, Lem. 1.3.2]). The projections $p^{X}_{k}\colon\lim_{\Pr^{R}}F_{*}\circ X\rightarrow X(k)$ admit left adjoint, denoted by $\gamma_{k}^{X}$. These left adjoints assemble to an equivalence
(2.21.1) $h_{X}\colon F_{K}(X)\xrightarrow{\sim}\lim_{k\in K^{\operatorname{op}},x_{*}}F(X_{i}).$
This is an equivalence in $\Pr^{L}$, where we view the RHS naturally as an object in $\Pr^{L}$. The inverse of $h_{X}$ is equivalent to $\mathop{\textup{colim}}_{k\in K}\textup{ins}^{X}_{k}\circ p_{k}^{X},$ where $\textup{ins}^{X}_{k}\colon F(X_{k})\rightarrow{\mathcal{M}}_{K}(X)$ denotes the canonical map.
Lemma 2.22.
Consider the setting of Lemma 2.20 (2). If $f\colon X\rightarrow Y$ is a morphism in $\operatorname{Fun}(K,{\mathcal{E}})$ such that each $f_{k}$ admits a left adjoint for all $k\in K$, then $f^{*}\coloneqq F(f)$ admits a left adjoint.
Proof.
This lemma is stated without proof in [BKV22, Prop. 5.1.8 (c)]. As far as we know, there is no written proof for this result in the literature, so we provide one.
By Lemma 2.20 (1), we see that $f_{*}\colon\lim_{K}F_{*}\circ X\rightarrow\lim_{K}F_{*}\circ Y$ admits a left adjoint $f^{*}$, that has to be compatible with the identification of $\lim_{K}F_{*}$ and $F_{K}$. We claim that the following diagram is commutative (up to homotopy)
${\lim_{K}F_{*}\circ Y}$${F(Y_{i})}$${\lim_{K}F_{*}\circ X}$${F(X)}$$\scriptstyle{p_{i}^{Y}}$$\scriptstyle{f^{*}}$$\scriptstyle{f_{i}^{*}}$$\scriptstyle{p_{i}^{X}}$
for all $i\in K$. This claim implies that $f^{*}$ commutes with limits and therefore admits a left adjoint $f_{\sharp}$.
Let us show the claim. For this, note that we have the following diagram with commutative squares (up to homotopy)
${F(Y_{j})}$${F_{K}(Y)}$${\lim_{K}F_{*}\circ Y}$${F(X_{j})}$${F_{K}(X)}$${\lim_{K}F_{*}\circ X}$$\scriptstyle{\textup{ins}_{Y}^{j}}$$\scriptstyle{f_{j}^{*}}$$\scriptstyle{f^{*}}$$\scriptstyle{h_{Y}}$$\scriptstyle{f^{*}}$$\scriptstyle{\textup{ins}_{X}^{j}}$$\scriptstyle{h_{X}}$
for any $j\in K$. First, we show that $p_{i}^{Y}\circ f^{*}$ commutes with colimits. Indeed, it is enough to see that $p_{i}^{Y}\circ f^{*}\circ h_{X}\circ\textup{ins}_{X}^{j}$ commutes with colimits, then is equivalent to the induced map $F_{K}(X)\rightarrow F(Y_{i})$ in $\Pr^{L}$, which necessarily commutes with colimits. By Lemma 2.20 (2), we have
$p_{i}^{X}\circ f^{*}\circ h_{Y}\circ\textup{ins}_{Y}^{j}\simeq p_{i}^{Y}\circ h_{X}\circ\textup{ins}_{X}^{j}\circ f_{j}^{*}\simeq\mathop{\textup{colim}}_{i\rightarrow k,j\rightarrow k}X_{ik*}\circ X_{jk}^{*}\circ f_{j}^{*}.$
By construction each of the functors $x_{ik*},x_{jk}^{*},f_{j}^{*}$ commutes with colimits and thus also the left hand side of the equivalence. Analogously, one can show that $f_{i}^{*}\circ p_{i}^{Y}$ commutes with colimits. Therefore, it is enough to show that
$p_{i}^{X}\circ f^{*}\circ h_{Y}\circ\textup{ins}_{Y}^{j}\simeq f_{i}^{*}\circ p_{i}^{Y}\circ h_{Y}\circ\textup{ins}_{Y}^{j}.$
The left hand side is equivalence to $\mathop{\textup{colim}}_{i\rightarrow k,j\rightarrow k}X_{ik*}\circ X_{jk}^{*}\circ f_{j}^{*}$ by the above and analogously, we see that the right hand side is equivalent to $\mathop{\textup{colim}}_{i\rightarrow k,j\rightarrow k}f_{i}^{*}\circ Y_{ik*}\circ Y_{jk}^{*}$. Since each square in $F(f)$ is right adjointable these colimits are in fact equivalent, proving the claim and therefore assertion. ∎
Lemma 2.23.
Consider the setting of Lemma 2.20 (2). Let $f$ be a morphism in $\operatorname{Fun}(K,{\mathcal{E}})$ such that $f_{k}$ admits a left adjoint for all $k\in K$. Let $f_{\sharp}$ denote the left adjoint of $f$ given by Lemma 2.22. Then the following square
${F(X_{k})}$${F_{K}(X)}$${F(Y_{k})}$${F_{K}(Y)}$$\scriptstyle{\textup{ins}_{X}^{k}}$$\scriptstyle{f_{k\sharp}}$$\scriptstyle{f_{\sharp}}$$\scriptstyle{\textup{ins}_{Y}^{k}}$
commutes (up to homotopy) for all $k\in K$.
Proof.
By construction of $f_{\sharp}$ the square
${F_{K}(X)}$${\lim_{K}F_{*}\circ X}$${F_{K}(Y)}$${\lim_{K}F_{*}\circ Y}$$\scriptstyle{h_{X}}$$\scriptstyle{f_{\sharp}}$$\scriptstyle{f_{\sharp}}$$\scriptstyle{h_{Y}}$
commutes (up to homotopy). Thus, we may check if the square in the lemma commutes after applying $h_{X}$. But the proof of Lemma 2.22 shows that the square
${\lim_{K}F_{*}\circ Y}$${F(Y_{i})}$${\lim_{K}F_{*}\circ X}$${F(X_{k})}$$\scriptstyle{p_{k}^{Y}}$$\scriptstyle{f^{*}}$$\scriptstyle{f_{k}^{*}}$$\scriptstyle{p_{k}^{X}}$
commutes (up to homotopy) for all $k\in K$. So in particular, we conclude by passing to left adjoints. ∎
Proof of Proposition 2.18.
Assertion (1) follows from Lemma 2.22.
We are left to show (2) and (3). Invoking Lemma 2.20 resp. Lemma 2.23 and the fact that each compact of ${\mathcal{M}}(X)$ resp. ${\mathcal{M}}(Y)$ comes from some level $i\in K$ (cf. Lemma 2.19), we can reduce the questions to questions on ${\mathcal{C}}$. But on ${\mathcal{C}}$ this follows from definition. ∎
Remark 2.24.
Let us consider the setting of Example 2.15 (2). By Proposition 2.18 we immediately see that the functor $\textup{DM}\colon(\textup{ProAlg}_{S}^{{\mathcal{I}}})^{\operatorname{op}}\rightarrow\textup{CAlg}(\Pr^{L}_{\omega})$ defines a pullback formalism with respect to the class of adjointable smooth morphisms.
2.2 Coefficient systems
In this short subsection, we want to show that our extended pullback formalisms allow us to define coefficient systems on $-{\mathcal{I}}$-algebraic stacks. We have to be a bit more precise, when formulating homotopy invariance $T$-stability and localization. For example $\mathbb{A}^{1}$-invariance makes only sense for cartesian morphisms, i.e. the $-{\mathcal{I}}$-algebraic version of $\mathbb{A}^{1}$ needs to be pulled back from $\mathbb{A}^{1}_{S}$. This will be made more precise in the Corollary 2.26.
We first start with the definition of a coefficient system over an arbitrary scheme $Q$ following Drew (cf. [Dre18]). We want to remark that we do only work with compactly generated. The main reason is that we have to use Proposition 2.18 that we only know in this case.
Definition 2.25 ([Dre18]).
A coefficient system is a compactly generated pullback formalism
${\mathcal{M}}^{*,\otimes}\colon(\textup{AffSch}^{\kappa}_{Q})^{\operatorname{op}}\rightarrow\textup{CAlg}(\textup{Pr}^{L,\otimes}_{\omega}),\quad f\mapsto f^{*}\quad$
satisfying the following conditions.
  1. (Loc)
    For any closed immersion $i\colon Z\hookrightarrow X$ in $\textup{Sch}_{Q}$ with open complement $j\colon U\hookrightarrow X$, the square
[…]
    is a pullback in $\Pr^{L}$.
  2. (Ho)
    For each $X\in\textup{Sch}_{Q}$ the *-pullback along the natural projection $p\colon\mathbb{A}^{1}_{X}\rightarrow X$ is fully faithful, i.e. the unit $\operatorname{id}\rightarrow p_{*}p^{*}$ is an equivalence.
  3. (St)
    For any smooth morphism $f\colon X\rightarrow Y$ in $\textup{Sch}_{Q}$ admitting a section $s$, the transformation $f_{\sharp}s_{*}$ is an equivalence.
Let ${\mathcal{M}}^{*\otimes}$ be a coefficient system. Then we define the class ${\mathcal{P}}_{{\mathcal{I}}}$ similarly to the definition in Proposition 2.18. To be more precise, ${\mathcal{P}}_{{\mathcal{I}}}$ consists of smooth adjointable morphisms $f$ in $\textup{ProAlg}^{{\mathcal{I}}}$. By Proposition 2.18 it is easy to see that ${\mathcal{M}}_{{\mathcal{I}}}$ defines a coefficient system with respect to the class ${\mathcal{P}}_{{\mathcal{I}}}$ in the following sense.
Corollary 2.26.
Let ${\mathcal{M}}^{*,\otimes}$ be a coefficient system that satisfies étale descent. Then ${\mathcal{M}}_{{\mathcal{I}}}$, that is the right Kan extension of ${\mathcal{M}}^{*,\otimes}$ along Yoneda to Artin $S$-stacks locally of finite type and then the composition of $\operatorname{Fun}(K,{\mathcal{M}}^{*\otimes})$ with the colimit functor, is a pullback formalism in the following sense.
  1. (PF)
    The functor ${\mathcal{M}}_{{\mathcal{I}}}$ defines a pullback formalism for the class ${\mathcal{P}}_{{\mathcal{I}}}$.
  2. (Loc)
    Let $i\colon(z,Z)\hookrightarrow(x,X)$ be cartesian closed immersion of $-{\mathcal{I}}$-algebraic stacks with open complement $j\colon(u,U)\rightarrow(x,X)$. Then there following diagram
[…]
    is a pullback in $\Pr^{L}$.
  3. (Ho)
    Let $(x,X)$ be a $-{\mathcal{I}}$-algebraic stack and let us define $(q,\mathbb{A}^{1}_{X})$ via the system $n\mapsto\mathbb{A}^{1}_{X_{i}}$. Then $p^{*}$ is fully faithful, where $p\colon(q,\mathbb{A}^{1}_{X})\rightarrow(x,X)$ denotes the projection.
  4. (St)
    For any smooth adjointable morphism $f\colon(x,X)\rightarrow(y,Y)$ of $-{\mathcal{I}}$-algebraic stacks, admitting a section $s$, the transformation $f_{\sharp}s_{*}$ is an equivalence.
Proof.
Claim (PF) follows from Proposition 2.18. Similar to the proof of (2) and (3) in Proposition 2.18, we can reduce to the case of Artin stacks via Lemma 2.20 and Lemma 2.19. For Artin stacks this follows immediately from descent. ∎
Remark 2.27.
The proof of Corollary 2.26 shows that if we restrict ourselves to pro-schemes locally of finite type over $S$, then we may drop the descent assumption on our coefficient system.
2.3 Six functor formalisms
In this section, we want to prove Theorem 5. For this, we need to extend $6$-functor formalisms to pro-algebraic stacks. We will freely use the formalism of Gaitsgory-Rozenblym, Liu-Zheng and Mann (cf. [GR17], [LZ12], [Man22]). We will use the results of the latter and for details refer to [Man22, §A.5]. We do not want to recall the main definitions and freely use the language of geometric setups and correspondences as in loc. cit..
Let $({\mathcal{C}},E)$ be a geometric setup, that is an $\infty$-category ${\mathcal{C}}$ and a class of morphisms $E$ in ${\mathcal{C}}$ that is closed under pullbacks, compositions and equivalences, such that ${\mathcal{C}}$ admits finite limits. Let $K$ a filtered $\infty$-category. Let ${\mathcal{D}}\colon\textup{Corr}({{\mathcal{C}}})_{E}\rightarrow\textup{Cat}_{\infty}$ be a $6$-functor formalism in the sense of [Man22, §A.5], where $\textup{Corr}({{\mathcal{C}}})_{E}$ is the $\infty$-category of correspondences in the sense of [LZ12].
We let ${\mathcal{D}}_{{\mathcal{C}}}\colon{\mathcal{C}}^{\operatorname{op}}\rightarrow\textup{CAlg}(\textup{Cat}_{\infty})$ denote the restriction of ${\mathcal{D}}$ under the inclusion ${\mathcal{C}}^{\operatorname{op}}\hookrightarrow\textup{Corr}({{\mathcal{C}}})_{E}$. We assume that ${\mathcal{D}}_{{\mathcal{C}}}$ factors through $\textup{CAlg}(\textup{Pr}^{L,\otimes}_{\omega})$. We will use a similar notation as in Section 2.1. We denote by ${\mathcal{D}}_{K}$ the composition
[…]
We define $E_{K}$ to be the class of those morphisms $f\colon X\rightarrow Y$ in $\operatorname{Fun}(K,{\mathcal{C}}^{\operatorname{op}})$ such that $f_{k}$ is in $E$ for all $k\in K$ and for all $k\rightarrow l\in K$ the square
[…]
is cartesian.
Proposition 2.28.
Assume that $K$ admits an initial object 00 and $E$ admits a suitable decomposition $J,F$ that satisfies the assumptions of [Man22, Prop. A.5.10]. Then ${\mathcal{D}}_{K}$ can be upgraded to a $6$-functor formalism
${\mathcal{D}}_{K}\colon\textup{Corr}({\operatorname{Fun}(K^{\operatorname{op}},{\mathcal{C}})})_{E_{K}}\rightarrow\textup{CAlg}(\textup{Pr}^{L,\otimes}_{\omega}).$
Further, if $f\in E$ is a morphism such that for all $k\in K$ we have $f_{k}\in F$ resp. $f(k)\in J$, then $f_{!}\simeq f_{*}$ resp. $f_{!}$ is left adjoint to $f^{*}$.
Proof.
We define $J_{K}$ resp. $F_{K}$ to be the class of those morphisms $f$ in $E_{K}$ such that $f(k)$ is in $J$ resp. $F$ for all $k\in K$. Then we claim that $J_{K},F_{K}$ is a suitable decomposition of $E$.
Indeed, since for a map $f$ in $E$ all squares of the associated diagram are pullback squares. Thus, all of the properties follow immediately (note that we need that $K$ admits an initial object for the decomposition property).
We will use Proposition 2.18 and the criterion in [Man22, Proposition. A.5.10] to prove this proposition.
By Proposition 2.18 the functor ${\mathcal{D}}_{K}$ defines a pullback formalism on $\operatorname{Fun}(K,{\mathcal{C}})$ with respect to $J_{K}$.
By construction, the pullback $f^{*}$ along any morphism $f\in\operatorname{Fun}(K,{\mathcal{C}})$ admits a right adjoint. We need to check that for morphisms in $F_{K}$ we have base change and projection formula. But similar to the proof of Proposition 2.18, we can use that any compact object comes from some compact on the $k$-th level, for some $k\in K$ (by mutas mutandis of Lemma 2.19). This argument shows that for any pullback square
${W}$${X}$${Y}$${Z}$$\scriptstyle{f^{\prime}}$$\scriptstyle{g^{\prime}}$$\scriptstyle{g}$$\scriptstyle{f}$
with $g\in J_{K}$ and $f\in F_{K}$ we have $g_{\sharp}f^{\prime}_{*}\simeq f_{*}g^{\prime}_{\sharp}$ completing the proof. ∎
The above proposition a priori only applies if the geometric setup has a suitable decomposition. For schemes this is usually not a problem as in practice we can decompose separated morphisms via compactifications into open immersions and proper maps. For algebraic stacks this only holds locally. Nevertheless, we can extend the functors by the base change formula to cartesian morphisms of $-{\mathcal{I}}$-algebraic stacks (cf. Proposition 2.29 below).
Proposition 2.29.
Let $f\colon X\rightarrow Y$ be in $E_{K}$. Then there exists an adjunction
${f_{!}\colon{\mathcal{D}}_{K}(X)}$${{\mathcal{D}}_{K}(Y)\colon f^{!}}$
satisfying base change and projection formula with respect to $f^{*}\coloneqq{\mathcal{D}}_{K}(f)$.
Proof.
Let $\widetilde{{\mathcal{D}}}$ denote the composition
[…]
Note that by design we have $\widetilde{{\mathcal{D}}}_{|\operatorname{Fun}(K,{\mathcal{C}}^{\operatorname{op}})}\simeq{\mathcal{D}}_{K}$. We will construct a functor $F\colon K\times\Delta^{1}\rightarrow\textup{Corr}({{\mathcal{C}}})_{E}$ that automatically yields a colimit preserving functor $f_{!}$ satisfying the necessary properties via $\widetilde{{\mathcal{D}}}$.
Let $\sigma\coloneqq(\sigma_{K},\sigma_{0,1})$ be an $n$-simplex in $K\times\Delta^{1}$. If $\sigma_{0,1}$ is a degeneration of $0\in\Delta^{1}$, then we define $F(\sigma)\coloneqq X(\sigma_{K})$ and if it is a degeneration of $1\in\Delta^{1}$, then we define $F(\sigma)\coloneqq Y(\sigma_{K})$.
Now assume that $\sigma_{0,1}$ is not a degeneration of a 00-simplex. Let $d$ denote the face map that selects the final vertex of an $n$-simplex. We will construct a correspondence $C_{X,Y}$ inside $\textup{Corr}({{\mathcal{C}}})_{E}$ as follows. As before, we let $f_{d(\sigma_{K})}$ be the morphism $f(d(\sigma_{K}),\langle 0,1\rangle)\colon X(d(\sigma_{K}))\rightarrow Y(d(\sigma_{K}))$ in ${\mathcal{C}}$. This defines a morphism $\alpha$ of correspondences
${X(d_{f}(\sigma_{K}))}$${X(d_{f}(\sigma_{K}))}$${Y(d_{f}(\sigma_{K})).}$$\scriptstyle{f_{d(\sigma_{K})}}$
We let $d^{c}$ denote the face map that forgets about the final vertex, i.e. $d^{c}$ is induced by the map $[n-1]\mapsto[n]$ inside $\Delta$ whose image omits the element $n$. Then $Y(d^{c}(\sigma_{K}))$ defines a $(n-1)$-simplex $\beta$ in $\textup{Corr}({{\mathcal{C}}})_{E}$ via the inclusion ${\mathcal{C}}^{\operatorname{op}}\hookrightarrow\textup{Corr}({{\mathcal{C}}})_{E}$. Informally, we may think of $\beta$ as the composition of $Y(d_{n-2,n-1}(\sigma_{K}))^{*}Y(d_{n-3,n-2}(\sigma_{K}))^{*}\dots Y(d_{0,1}(\sigma_{K}))^{*}$, where the $d_{i,j}$ denotes the edge connecting the $i$-th and $j$-th vertex inside $d^{c}(\sigma_{K})$. The morphism $\alpha$ defines an edge to the initial vertex of $Y(d^{c}(\sigma_{K}))$. The extension of this edge yields an $n$-simplex, which we may describe as the composition $\beta\circ\alpha$. We now set $C_{X,Y}\coloneqq\beta\circ\alpha$. Since by design $f$ is cartesian this construction is functorial and yields the desired functor $F\colon\sigma\mapsto C_{X,Y}$. The composition of $F$ with ${\mathcal{D}}$ corresponds to maps $f_{k!}\colon X(k)\rightarrow Y(k)$ compatible with the $X_{kj}^{*}$ resp. $Y_{kj}^{*}$ for all $k\rightarrow j\in K$ (also in higher dimensions by construction). In particular, taking colimits inside $\Pr^{L}$ yields a colimit preserving functor $f_{!}\colon{\mathcal{D}}_{K}(X)\rightarrow{\mathcal{D}}_{K}(Y)$, as desired.
The rest follows from Lemma 2.19 similar to the proof of Proposition 2.18.
Remark 2.30.
Note that it is clear from the proof of Proposition 2.29 that if ${\mathcal{D}}^{!}_{|{\mathcal{C}}_{E}}$ commutes with colimits that also $f^{!}\coloneqq{\mathcal{D}}_{K}^{!}(f)$ commutes with colimits for all $f\in E_{K}$ as ${\mathcal{D}}_{K}(X)$ is compactly generated for all $X\in\operatorname{Fun}(K,{\mathcal{C}})$ by Lemma 2.19.
Proposition 2.31.
Let ${\mathcal{I}}\simeq\mathbb{N}_{0}$, the $\infty$-category associated to the poset $\mathbb{N}_{0}$ seen as a $1$-category. Let $(x,X)$ be a monoidal $-{\mathcal{I}}$-algebraic stack, then
$\textup{DM}(x,X)\simeq\mathop{\textup{colim}}_{i\in{\mathcal{I}},x^{!}}\textup{DM}^{!}(X_{i}).$
Proof.
By descent, we see that for each $i\in{\mathcal{I}}$ we have $\tilde{p}_{i}\colon\textup{DM}^{!}(X_{i})\simeq\textup{DM}^{*}(X_{i})$ as $X_{i}$ admits a smooth covering by a scheme. We will construct equivalences
$p_{i}\colon\textup{DM}^{!}(X_{i})\simeq\textup{DM}^{*}(X_{i})$
out of $\tilde{p}_{i}$ that are compatible with the presentation as a colimit. This is easily achieved by twisting with the Thom-motives associated to the smooth maps $x_{0i}$. But let us be more precise.
Let us define $p_{i}\coloneqq\tilde{p}_{i}\otimes\textup{Th}(\Omega_{x_{0i}})1_{X_{i}}$, where Th denotes the Thom-motive [^1] Let $p\colon V\rightarrow X_{i}$ be a vector bundle with zero section $s$, then $\textup{Th}(V)\coloneqq p_{\sharp}s_{*}$. associated to $\mathbb{V}(\Omega_{x_{0i}})\rightarrow X_{i}$. Then $p_{i}$ is still an equivalence of $\textup{DM}^{!}(X_{i})$ and $\textup{DM}^{*}(X_{i})$. By compatibility with the purity equivalence with composition (cf. [CD19, Rem. 2.4.52]), we claim that this is also compatible with the presentation of $\textup{DM}(x,X)$ as a colimit.
Indeed, the functor $\textup{Spine}[\mathbb{N}_{0}]\rightarrow\mathbb{N}_{0}$ is inner anodyne [Lur21, 03HK], where $\textup{Spine}[\mathbb{N}_{0}]$ is the simplicial subset of $\mathbb{N}_{0}$ of finite vertices that are joined by edges (or more informally, we forget about all $n\geq 2$ simplices in $\mathbb{N}_{0}$). Moreover, by loc. cit., we see that the restriction $\operatorname{Fun}(\mathbb{N}_{0},\textup{DGCat}_{\textup{cont}})\rightarrow\operatorname{Fun}(\textup{Spine}[\mathbb{N}_{0}],\textup{DGCat}_{\textup{cont}})$ is a trivial Kan fibration. Thus, we can lift the map that we constructed to a map of diagrams indexed by $\mathbb{N}_{0}$ proving the claim. ∎
Remark 2.32.
Under the assumption of Proposition 2.31, we can use Proposition 2.18 and similar methods as in the proof to conclude that ${\mathcal{D}}_{{\mathcal{I}}}$ is part of a six functor formalism with respect to the class $\tilde{E}_{{\mathcal{I}}}$, where $\tilde{E}_{{\mathcal{I}}}$ is the class of all morphisms $f$ such that each square is right adjointable and $f_{i}\in E$ for all $i\in{\mathcal{I}}$. To be more precise, we get the existence of $f_{!}$ and $f^{!}$ for morphisms in $\tilde{E}_{{\mathcal{I}}}$ satisfying base change and projection formula.
Remark 2.33.
To generalize Proposition 2.31 it is enough to know that the level-wise equivalences $p_{i}$ in the proof yield an equivalence of diagrams. While on the homotopy categorical level this is clear, we were not able to write a map down for higher simplices.
Proof of Theorem 5.
By Ayoub’s thesis resp. Cisinski-Deglise (cf. [Ayo07] resp. [CD19]) there is a full motivic six functor formalism on Beilinson motives for finite type $S$-schemes. Further, by Robalo’s thesis [Rob15, §2.4], we have that the $6$-functor formalism exists in the sense of [LZ12]. Note that DM also satisfies étale descent. Thus, the proof of the theorem boils down to Proposition 2.31, Proposition 2.29, Remark 2.32, Lemma 2.22, Lemma 2.20 and Lemma 2.19, similar to the proof of Proposition 2.18. ∎
Remark 2.34.
The proof of Theorem 5 can be generalized to other settings such as ind-schemes. To be more specific, using the result of Richarz and Scholbach on rational motives of ind-schemes [RS20, Thm. 2.4.2], we can replace algebraic stacks of finite type, by ind-schemes of ind-finite type and modify DM as follows.
If $(x,X)$ is a $-{\mathcal{I}}$-system of ind-schemes of ind-finite type (let us call such systems $-{\mathcal{I}}$-schemes), then we define $\textup{DM}(x,X)\coloneqq\mathop{\textup{colim}}_{i\in{\mathcal{I}},x^{!}}\textup{DM}^{!}(X_{i})$, note that we use !!!-pullbacks as transition maps.
From now on assume that the transition maps of $-{\mathcal{I}}$-schemes are schematic smooth. Then $\textup{DM}(x,X)$ admits a $\otimes$-product and internal $\operatorname{Hom}$. The existence of the $\otimes$-unit fails in general (cf. loc. cit.). Let $f\colon(x,X)\rightarrow(y,Y)$ be morphism of $-{\mathcal{I}}$-schemes (we define this similarly to Definition 2.2). Assume each $f_{i}$ satisfies the assumptions of [RS20, Thm 2.4.2] for the existence of $f_{i}^{*}$. Then there exists an adjunction $f^{*}\dashv f_{*}$. In general there exists an adjunction $f_{!}\dashv f^{!}$, where we use that $(x,X)$ and $(y,Y)$ are required to be schematic smooth (in this way the right adjoint of $x^{!}$ is up to twist and shift given by $x_{*}$). All of these functors satisfy the properties of Theorem 5 using the same proof.
3 Rational motivic cohomology on pro-algebraic stacks
Before we continue with the definition of motivic cohomology, we want to insert some consequences of Theorem 5 and show that $\textup{DM}(x,X)\simeq\textup{DM}(X)$ for classical monoidal $-{\mathcal{I}}$-algebraic stacks.
Proposition 3.1.
Let $(x,X)$ be a classical monoidal $-{\mathcal{I}}$-algebraic stack. Then we have an equivalence
$\textup{DM}(X)\simeq\textup{DM}(x,X)$
of DG-categories.
Proof.
By assumption $X_{0}$ admits a smooth cover by affine finite type $S$-schemes $U_{0,i}=\operatorname{Spec}(A_{i})$, for some set $I$ and $i\in{\mathcal{I}}$. As $X_{i}\rightarrow X_{0}$ is affine, the pullback $U_{n,i}\coloneqq X_{i}\times_{X_{0}}U_{0,i}$ is also a smooth cover by affines. The same holds after passage to limits, i.e. for $U_{\infty,i}\coloneqq X\times_{X_{0}}U_{0,i}$ the projection $\coprod_{i\in{\mathcal{I}}}U_{\infty,i}\rightarrow X$ is an effective epimorphism. In particular, as DM satisfies $h$-descent (cf. [RS20, Thm. 2.2.16]), we have $\lim_{\Delta}\textup{DM}(\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{\infty,i}/X))\simeq\textup{DM}(X)$. By construction, each of the $U_{\infty,i}$ is equivalent to $\lim_{i\in{\mathcal{I}}}U_{n,i}$. As the $U_{n,i}$ are affine of finite type over $S$, we have
$\textup{DM}(U_{\infty,i})\simeq\textup{DM}(\lim_{i\in{\mathcal{I}}}U_{n,i})\simeq\mathop{\textup{colim}}_{i\in{\mathcal{I}},x^{*}}\textup{DM}(U_{n,i})\simeq\lim_{i\in{\mathcal{I}}^{\operatorname{op}},x_{*}}\textup{DM}(U_{n,i})$
(cf. Lemma 2.20). This yields the equivalence
$\textup{DM}(\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{\infty,i}/X))\simeq\lim_{i\in{\mathcal{I}},x_{*}}\textup{DM}(\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{n,i}/X_{i})).$
Putting all of this together, we claim that this yields
$\textup{DM}(X)\simeq\lim_{\Delta}\textup{DM}(\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{\infty,i}/X))$ $\simeq\lim_{i\in{\mathcal{I}}^{\operatorname{op}},x_{*}}\lim_{\Delta}\textup{DM}(\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{n,i}/X_{i}))$
$\simeq\lim_{i\in{\mathcal{I}}^{\operatorname{op}},x_{*}}\textup{DM}(X_{i})\simeq\mathop{\textup{colim}}_{i\in{\mathcal{I}},x^{*}}\textup{DM}(X_{i})$
as desired.
Indeed, the only thing to check is the second equivalence. For this it is enough to see that $x_{ij}^{*}$ induces a map of simplicial objects
$\textup{DM}(\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{n,i}/X_{i}))\rightarrow\textup{DM}(\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{n-1,i}/X_{j})).$
But this follows immediately from smooth base change, as
$\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{n,i}/X_{i})\simeq\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{n-1,i}/X_{j})\times_{X_{j}}X_{i}.$
Remark 3.2 (Underling motive of pro-algebraic stacks).
Let us show how to compute the underlying motive in $\textup{DM}(S)$ of a monoidal strict $-{\mathcal{I}}$-algebraic stack.
Let $(x,X)$ be a monoidal strict $-{\mathcal{I}}$-algebraic stack. We have seen in Lemma 2.20 that there is a map $p_{0}\colon\textup{DM}(x,X)\rightarrow\textup{DM}(X_{0})$ such that the identity on $\textup{DM}(X_{0})$ factors as
$\textup{DM}(X_{0})\xrightarrow{\textup{ins}_{0}}\textup{DM}(x,X)\xrightarrow{p_{0}}\textup{DM}(X_{0}),$
where $\textup{ins}_{0}$ denotes the natural inclusion.
Assume that ${\mathcal{I}}\simeq\mathbb{N}_{0}$, then by Theorem 5 the two morphisms above correspond to the adjunction
${c_{0*}\colon\textup{DM}(x,X)}$${\textup{DM}(X_{0})\colon c_{0}^{*}}$
where $c_{0}\colon(x,X)\rightarrow X_{0}$ denotes the projection, where we see $X_{0}$ as a constant $-{\mathcal{I}}$-algebraic stack. In particular, since $(x,X)$ is monoidal, both $c_{0}^{!}$ and $c_{0}^{*}$ are fully faithful.
Let us denote by $f\colon X\rightarrow S$ and $f_{0}\colon X_{0}\rightarrow S$ the structure maps and let $M\in\textup{DM}(S)$. The above shows that $f_{!}f^{!}M\simeq f_{0!}f_{0}^{!}M$.
Remark 3.3 (Base change for pro-algebraic stacks).
Let $f\colon(x,X)\rightarrow(y,Y)$ be a smooth morphism monoidal $-{\mathcal{I}}$-algebraic stacks. Assume that $f$ is either adjointable and ${\mathcal{I}}\simeq\mathbb{N}_{0}$ or that $f$ is cartesian and ${\mathcal{I}}$ has an initial object 00.
Our computations have shown that we have $f_{*}\textup{ins}_{0}^{X}\simeq\textup{ins}_{0}^{Y}f_{0*}$, where $\textup{ins}_{0}^{X}$ denotes the natural functor $\textup{DM}(X_{0})\rightarrow\textup{DM}(x,X)$ (similarly for $Y$). This can be see as a form of non-finite type base change equivalence of the square
${(x,X)}$${X_{0}}$${(y,Y)}$${Y_{0},}$
which is cartesian if $f$ is cartesian.
For example this yields a *-base change formula for pullback along the generic point of an integral scheme (cf. Example 3.13 below). This is not true in rational motives, if the inclusion of the generic point may not be of finite type.
Remark 3.4.
Let $f\colon(x,X)\rightarrow(y,Y)$ be a smooth morphism monoidal $-{\mathcal{I}}$-algebraic stacks. For $M\in\textup{DM}(y,Y)$, we can compute the "motivic global sections with values in $M$" directly. To be more precise, let us note that since $f_{*}f^{*}$ is colimit preserving, its values are determined by its restriction to compacts. As any compact in $\textup{DM}(y,Y)$ comes from a compact in some $\textup{DM}(Y_{i})$ (cf. Lemma 2.19), we see that
$f_{*}f^{*}M\simeq\mathop{\textup{colim}}_{\textup{ins}_{i}M_{i}\rightarrow M}\mathop{\textup{colim}}_{i\rightarrow k,j\rightarrow k}\textup{ins}_{i}^{Y}f_{i*}f^{*}_{i}y_{ik*}y_{ij}^{*}M_{i},$
where the colimit is over all $i\in{\mathcal{I}}$ and all compacts $M_{i}\in\textup{DM}(Y_{i})^{c}$. A similar formula can also be given for $f_{!}f^{!}$.
In particular, if for example $(y,Y)\cong S$, i.e. the constant diagram associated to $S$, and $(x,X)$ is strict, we have
$f_{*}f^{*}1_{S}\simeq f_{0*}f_{0}^{*}1_{S}$
(cf. Remark 3.2).
3.1 Motivic cohomology
In this section, we want to define and highlight some properties of motivic cohomology in our setting.
In the following, we fix a morphism $f\colon(x,X)\rightarrow(y,Y)$ of monoidal $-{\mathcal{I}}$-algebraic stacks over $S$. Let $h\colon(x,X)\rightarrow S$ and $g\colon(y,Y)\rightarrow S$ be the morphism induced by the structure morphisms, where we view $S$ as a constant $-{\mathcal{I}}$-algebraic stack. We further assume that either $f$ is cartesian or that $f$ is adjointable and ${\mathcal{I}}\simeq\mathbb{N}_{0}$.
We fix the notation as in Remark 2.21, i.e. we have an equivalence
$h_{X}\colon\textup{DM}(x,X)\rightarrow\lim_{i\in{\mathcal{I}},x_{*}}\textup{DM}(X_{i}),$
we denote for any $i\in{\mathcal{I}}$ by
$\textup{ins}_{i}^{X}\colon\textup{DM}(X_{i})\rightarrow\textup{DM}(x,X)\textup{ resp. }p^{X}_{i}\colon\lim_{i\in{\mathcal{I}},x_{*}}\textup{DM}(X_{i})\rightarrow\textup{DM}(X_{i})$
the canonical inclusion resp. projections and we set $\gamma_{i}^{X}\coloneqq h_{X}\circ\textup{ins}_{i}^{X}$ (analogously for $(y,Y)$).
By Theorem 5, we have adjunctions $f^{*}\dashv f_{*}$ and $f_{!}\dashv f^{!}$ between $\textup{DM}(x,X)$ and $\textup{DM}(y,Y)$ satisfying all the properties stated in the theorem. Note that even though $(x,X)$ and $(y,Y)$ are not assumed to be strict, we still have adjunctions $h^{*}\dashv h_{*}$ and $g^{*}\dashv g_{*}$ by Lemma 2.22.
Definition 3.5.
Let $f\colon(x,X)\rightarrow(y,Y)$ and $g\colon(y,Y)\rightarrow S$ be as above. Then we define the relative motivic cohomology of $(x,X)$ with coefficients in $M$ as
$R\Gamma(X,M)\coloneqq\underline{\operatorname{Hom}}_{{\mathcal{D}}(\mathbb{Q})}(f_{!}f^{!}1_{(x,X)},M).$
Further, we define for $n,m\in\mathbb{Z}$ the absolute motivic cohomology of $(y,Y)$ in degree $(n,m)$ by
$H^{n,m}(Y,\mathbb{Q})\coloneqq\pi_{0}\operatorname{Hom}_{\textup{DM}(y,Y)}(1_{(y,Y)},1_{(y,Y)}(n)[m]).$
Remark 3.6.
Let us remark that the absolute motivic cohomology of $(y,Y)$ can also be computed relative to $S$ in the following sense
$H^{n,m}(Y,\mathbb{Q})\cong\operatorname{Hom}_{h\textup{DM}(S)}(1_{S},g_{*}g^{*}1_{S}(n)[m]).$
Further, if $(y,Y)$ is strict and $(y,Y)\rightarrow S$ is a smooth, then
$H^{n,m}(Y,\mathbb{Q})\cong\pi_{m}\operatorname{Hom}_{\textup{DM}(S)}(g_{\sharp}g^{*}1_{S},1_{S}(n)).$
Lemma 3.7.
Under the notation above, we have
$g_{*}1_{(y,Y)}\simeq\mathop{\textup{colim}}_{j\in{\mathcal{I}}}g_{j*}1_{Y_{j}}.$
Proof.
By construction of $g^{*}$ in Lemma 2.22, we have $\mathop{\textup{colim}}_{i\in{\mathcal{I}}}\textup{ins}_{i}^{Y}g^{*}_{i}$. Using adjunctions this shows that
$g_{*}\simeq\lim_{i}g_{i*}p^{Y}_{i}h_{Y}.$
By construction, we have $\textup{ins}^{Y}_{0}1_{Y_{0}}\simeq 1_{(y,Y)}$. Thus,
$g_{*}1_{(y,Y)}\simeq\lim_{i}g_{i*}p^{Y}_{i}\gamma_{0}^{Y}1_{Y_{0}}$ $\simeq\lim_{i}\mathop{\textup{colim}}_{i\rightarrow j}g_{i*}y_{ij*}y^{*}_{0j}1_{Y_{0}}$
$\simeq\lim_{i}\mathop{\textup{colim}}_{i\rightarrow j}g_{j*}1_{Y_{j}}$
$\simeq\lim_{i}\mathop{\textup{colim}}_{j\in{\mathcal{I}}}g_{j*}1_{Y_{j}}$
$\simeq\mathop{\textup{colim}}_{j\in{\mathcal{I}}}g_{j*}1_{Y_{j}}.$
Remark 3.8.
Under the notation above, Lemma 3.7 immediately shows
$H^{n,m}(Y,\mathbb{Q})\cong\mathop{\textup{colim}}_{i\in{\mathcal{I}}}A^{n}(Y_{i},2n-m)_{\mathbb{Q}}.$
We will see below, that if $(y,Y)$ is classical and $Y\rightarrow S$ is of finite type, then this colimit presentation of the absolute motivic cohomology of $Y$ agrees with the classical motivic cohomology of $Y$. In this setting, we can also look at the functor induced by the usual *-pushforward on Artin-stacks $\textup{DM}(y,Y)\simeq\textup{DM}(Y)\rightarrow\textup{DM}(S)$. This agrees by construction with $g_{*}\colon\textup{DM}(y,Y)\rightarrow\textup{DM}(S)$.
Further, in Example 3.12 and 3.13, we can use this computation for the absolute motivic cohomology, even in the non-finite type case (see the examples for an elaborate discussion).
Lemma 3.9.
Assume $(y,Y)$ is classical such that $g\colon Y\rightarrow S$ is locally of finite type. Let us denote by $\alpha\colon\textup{DM}(y,Y)\xrightarrow{\sim}\textup{DM}(Y)$ the equivalence induced by Proposition 3.1. Let us denote by $\tilde{g}\colon\textup{DM}(Y)\rightarrow\textup{DM}(S)$ the usual *-pushforward. Then we have
$g_{*}1_{(y,Y)}\simeq\tilde{g}\alpha 1_{(y,Y)}.$
Proof.
By construction, the last right hand side of Lemma 3.7 is equivalent to $\tilde{g}\alpha 1_{(y,Y)}$. ∎
Proposition 3.10.
Assume $(y,Y)$ is classical such that $g\colon Y\rightarrow S$ is locally of finite type. Then
$H^{n,m}(Y,\mathbb{Q})\simeq\operatorname{Hom}_{h\textup{DM}(Y)}(1_{Y},1_{Y}(n)[m]).$
Proof.
Let $\tilde{Y}_{0}\rightarrow Y_{0}$ be a smooth cover of $Y_{0}$ by a scheme locally of finite type over $S$. Let $\tilde{Y}_{i}$ denote the base change to $Y_{i}$ and $\tilde{Y}$ its limit. Then $\tilde{Y}$ is representable by a scheme and yields a smooth cover of $Y$. As $Y$ is locally of finite type, so is $\tilde{Y}$. For every $i\in{\mathcal{I}}$, we can compute
$g_{i*}1_{Y_{i}}\simeq\mathop{\textup{colim}}_{\Delta}g_{\check{C}(\tilde{Y}_{i}/Y_{i})_{\bullet}*}1_{\check{C}(\tilde{Y}_{i}/Y_{i})_{\bullet}}$
by descent, where $g_{\check{C}(\tilde{Y}_{i}/Y_{i})_{\bullet}}\colon\check{C}(\tilde{Y}_{i}/Y_{i})_{\bullet}\rightarrow S$ denotes the projection. In particular, we can write
$f_{*}1_{(y,Y)}\simeq\mathop{\textup{colim}}_{j\in{\mathcal{I}}}g_{j*}1_{Y_{j}}\simeq\mathop{\textup{colim}}_{\Delta}\mathop{\textup{colim}}_{j\in{\mathcal{I}}}g_{\check{C}(\tilde{Y}_{i}/Y_{i})_{\bullet}*}1_{\check{C}(\tilde{Y}_{i}/Y_{i})_{\bullet}}$
(cf. Lemma 3.7). Therefore, by continuity of DM (cf. [CD19, Thm. 14.3.1]) and descent, we have
$H^{n,m}(Y,\mathbb{Q})$ $\cong\mathop{\textup{colim}}_{\Delta}\mathop{\textup{colim}}_{i\in{\mathcal{I}}}\operatorname{Hom}_{h\textup{DM}(\check{C}(\tilde{Y}_{i}/Y_{i})_{\bullet})}(1_{\check{C}(\tilde{Y}_{i}/Y_{i})_{\bullet}},1_{\check{C}(\tilde{Y}_{i}/Y_{i})_{\bullet}}(n)[m])$
$\cong\mathop{\textup{colim}}_{\Delta}\operatorname{Hom}_{h\textup{DM}(\check{C}(\tilde{Y}/Y)_{\bullet})}(1_{\check{C}(\tilde{Y}/Y)_{\bullet}},1_{\check{C}(\tilde{Y}/Y)_{\bullet}}(n)[m])$
$\cong\operatorname{Hom}_{h\textup{DM}(Y)}(1_{Y},1_{Y}(n)[m]).$
Remark 3.11.
Let us remark that continuity of DM (cf. [CD19, Thm. 14.3.1]) shows that if $(y,Y)$ is a pro system of finite type $S$-schemes such that $Y$ is representable by a finite type $S$-scheme, we have again $H^{n,m}(Y,\mathbb{Q})\cong\operatorname{Hom}_{h\textup{DM}(Y)}(1_{Y},1_{Y}(n)[m])$.
Example 3.12.
[…]
$H^{n,m}(Z_{L},\mathbb{Q})\cong\mathop{\textup{colim}}_{E\in{\mathcal{F}}^{L}_{K}}A^{n}(Z_{E},2n-m)_{\mathbb{Q}}\cong A^{n}(Z_{L},2n-m)_{\mathbb{Q}},$
where the second isomorphism follows from direct computations on the level of cycles (cf. [Bro16, Lem. 1.4.6 (i)]).
Example 3.13.
Let $Z\rightarrow S$ be an integral $S$-scheme with function field $K$. Let us consider the family $(U)_{U\in{\mathcal{U}}_{K}}$ of all affine open subschemes of $Z$. If we denote the respective inclusions by $\iota_{UU^{\prime}}\colon U^{\prime}\hookrightarrow U$, then this family assembles to a classical monoidal pro-${\mathcal{U}}_{K}$-algebraic stack $(\iota,\operatorname{Spec}(K))$. Let $T\rightarrow Z$ be a flat morphism of $S$-schemes. Then base change of $T$ along the system $(\iota,\operatorname{Spec}(K))$ induces a classical monoidal pro-${\mathcal{U}}_{K}$-algebraic stack $(\iota_{T},T_{K})$. Using Remark 3.6, we have
$H^{n,m}(T_{K},\mathbb{Q})\cong\mathop{\textup{colim}}_{U}A^{n}(T_{U},2n-m)_{\mathbb{Q}}\cong A^{n}(T_{K},2n-m)_{\mathbb{Q}},$
where the second isomorphism follows from direct computations on the level of cycles (cf. [Bro16, Lem. 1.4.6 (ii)]).
4 Applications
[…]
Next, we will apply our machinery to the stack of displays Disp (cf. Section 4.2). The stack Disp can be expressed as a strict monoidal $-\mathbb{N}$-algebraic stack $(\tau,\textup{Disp})$. Thus, the underling motive of $(\tau,\textup{Disp})$ is equivalent to the motive of its $1$-truncation. In particular, this allows us to compute the absolute motivic cohomology of Disp and by comparison also of the stack of Barsotti-Tate groups. Further, by earlier computation of the author, we see that the motive of Disp inside $\textup{DM}(\mathbb{F}_{p})$ is Tate.
4.1 Action of the absolute Galois group on the motivic homology spectrum
[…]
$f_{\bullet}\colon(x,{\mathfrak{X}}_{L})\rightarrow(q,L)$
Let ${\mathcal{G}}^{L}_{K}$ denote the filtered subcategory of all finite Galois extensions of $k$. Then ${\mathcal{G}}^{L}_{k}\subseteq{\mathcal{F}}^{L}_{K}$ is cofinal. In particular, we see that
$\textup{DM}(x,{\mathfrak{X}}_{L})\simeq\mathop{\textup{colim}}_{{\mathcal{G}}^{L}_{K},x^{*}}\textup{DM}(x,{\mathfrak{X}}_{L}).$
Thus, to understand the action of $\operatorname{Gal}(L/k)$ on $f_{L*}1_{{\mathfrak{X}}_{L}}$, we may restrict $(x,{\mathfrak{X}}_{L})$ and $(q,L)$ to ${\mathcal{G}}^{L}_{k}$ and work with their underlying pro-${\mathcal{G}}^{L}_{k}$-algebraic stacks. By abuse of notation, we will keep the notation of $(x,{\mathfrak{X}}_{L})$ and $(q,L)$.
We naturally get a map $g\colon(q,\operatorname{Spec}(L))\rightarrow\operatorname{Spec}(k)$, where we view $\operatorname{Spec}(k)$ as a constant diagram. We denote the composition $g\circ f_{\bullet}$ by $h_{\bullet}$.
Remark 4.1.
Note that Proposition 3.1 shows that
$\textup{DM}(x,{\mathfrak{X}}_{L})\simeq\textup{DM}({\mathfrak{X}}_{L})\textup{and }\textup{DM}(q,L)\simeq\textup{DM}(L).$
We will see in the next lemma that our formalism enables us to compute the motive $f_{L*}1_{{\mathfrak{X}}_{L}}$ by "pulling back" $f_{*}1_{{\mathfrak{X}}}$ along the natural map $\textup{DM}(k)\rightarrow\textup{DM}(q,L)$.
Assume that $L/k$ is a finite extension. Then the smooth base change immediately yields
$f_{L*}1_{{\mathfrak{X}}}\simeq f_{L*}x_{L/k}^{*}1_{{\mathfrak{X}}}\simeq q_{L/k}^{*}f_{*}1_{{\mathfrak{X}}}.$
As expected this also holds by the lemma below, if $L/k$ is not finite. If $L/k$ is not finite this does not hold as there is no base change formalism that shows this result. In our formalism we have an analog of a smooth base change in this case as in Remark 3.3.
Next, we claim that $\operatorname{Gal}(L/k)$ acts on $h_{*}1_{{\mathfrak{X}}_{L}}$. To see this, it is enough to construct actions of $\operatorname{Gal}(E/k)$ on $h_{*}1_{{\mathfrak{X}}_{L}}$ for any finite Galois extension $E/k$ that is compatible with the presentation of $\operatorname{Gal}(L/k)$ as a limit of such.
Remark 4.2.
Let $E/k$ be a finite Galois extension and $\varphi\in\operatorname{Aut}_{k}(E)$. And $M\in\textup{DM}(q,\operatorname{Spec}(L))$. Then $\varphi$ induces an automorphism
$g_{*}M\rightarrow g_{*}M$
via the following.
Let ${\mathcal{G}}_{E}^{L}\subseteq{\mathcal{G}}_{k}^{L}$ denote the subcategory of finite Galois extensions over $E$. Then we can restrict $q$ along this (filtered) subcategory and we denote the induced pro-${\mathcal{G}}_{E}^{L}$-algebraic stack by $(q_{|E},\operatorname{Spec}(L))$. The $k$-automorphism $\varphi$ induces via base change a cartesian automorphism $\varphi\colon(q_{|E},\operatorname{Spec}(L))\rightarrow(q_{|E},\operatorname{Spec}(L))$. The *-pushforward along $\varphi$ yields a functor
$\varphi_{*}\colon\textup{DM}(q,\operatorname{Spec}(L))\rightarrow\textup{DM}(q,\operatorname{Spec}(L)).$
By construction $g_{*}\varphi_{*}\simeq g_{*}$, yielding the desired endomorphism above.
This endomorphism by construction induces an action of $\operatorname{Gal}(L/k)$ on $g_{*}M$, i.e. a map $B\operatorname{Gal}(L/k)\rightarrow\textup{DM}(k)$.
Remark 4.3.
Combining the above, we have
$\mathop{\textup{colim}}_{E\in{\mathcal{G}}_{k}^{L}}h_{*}1_{{\mathfrak{X}}}^{\operatorname{Gal}(E/k)}\simeq\mathop{\textup{colim}}_{E\in{\mathcal{G}}_{k}^{L}}\mathop{\textup{colim}}_{F\in{\mathcal{G}}_{k}^{L}}q_{F/k*}f_{F*}1_{{\mathfrak{X}}_{F}}^{\operatorname{Gal}(E/k)}\simeq f_{*}1_{{\mathfrak{X}}}.$
(cf. [Ayo07, Lem. 2.1.166]). In particular, for ${\mathfrak{X}}=[X/G]$ with $X$ smooth $k$-scheme and $G$ linear algebraic $k$-group, we have
$A^{n}_{G}(X,m)_{\mathbb{Q}}\simeq A^{n}_{G_{L}}(X_{L},m)^{\operatorname{Gal}(L/k)}_{\mathbb{Q}}$
by Example 3.12. This recovers the computation on the level of cycles purely formally (cf. [Bro18, Lem. 1.3.6]).
There exists a notion of continuous homotopy fixed points on so-called discrete $G$-spectra (cf. [BD10]). The object $\mathop{\textup{colim}}_{E\in{\mathcal{G}}_{k}^{L}}h_{*}1_{{\mathfrak{X}}}^{\operatorname{Gal}(E/k)}$ can be seen as an analog of continuous homotopy fixed points under the action of the absolute Galois group. By construction the action of $\operatorname{Gal}(L/k)$ is stabilized on an open and closed normal subgroup after passage to a large enough field extension.
4.2 The motive of the stack of Displays
Let us fix a prime $p>0$ and assume from now on that $S=\operatorname{Spec}(\mathbb{F}_{p})$. We will denote the Frobenius with $\sigma$. In the following section, we want to define the $-\mathbb{N}$-algebraic stack of (truncated) Displays over $S$, after the construction of Lau (cf. [Lau13]) and show that its underlying motive in $\textup{DM}(S)$ is Artin-Tate. We do not want to go into the detail of the construction of (truncated) Displays and its relation to the stack of Barsotti-Tate groups, but rather give an equivalent definition following [Bro18].
Let us consider the functor $\operatorname{Spec}(R)\mapsto\operatorname{GL}_{h}(W_{n}(R))$ for any affine $S$-scheme $\operatorname{Spec}(R)$, where $W_{n}$ denotes the ring of $n$-truncated Witt-vectors. This functor is represented by an open subscheme of $\mathbb{A}^{nh^{2}}$, which we denote by $X_{n}^{h}$. Let $G^{h,d}_{n}(R)$ denote the group of invertible $h\times h$ matricies
$\begin{pmatrix}A&B\\ C&D\end{pmatrix}$
with $A\in\operatorname{GL}_{h-d}(W_{n}(R))$. Let us denote by $G^{h,d}_{n}$ the group scheme representing this functor. Then $G^{h,d}_{n}$ acts on $X^{h}_{n}$ via $M\cdot x\coloneqq Mx\sigma^{\prime}(M)^{-1}$, where
$\sigma^{\prime}(M)\coloneqq\begin{pmatrix}\sigma A&p\sigma B\\ \sigma C&\sigma D\end{pmatrix}.$
Definition 4.4.
For $0\leq d\leq h$ and $n\in\mathbb{N}$, we define the stack of $n$-truncated displays of dimension $d$ and height $n$ to be
$\textup{Disp}_{n}^{h,d}\coloneqq\left[X^{h}_{n}/G^{h,d}_{n}\right].$
Further, we define the stack of $n$-truncated displays as
$\textup{Disp}_{n}\coloneqq\coprod_{0\leq d\leq h}\textup{Disp}_{n}^{h,d}.$
Remark 4.5.
Let us remark that this is not the usual definition of the stack of truncated displays but rather an equivalent one (cf. [Bro18, Thm. 2.1.3]).
Notation 4.6.
For each $n\in\mathbb{N}$, there exists a truncation map $\tau_{n}\colon\textup{Disp}_{n+1}\rightarrow\textup{Disp}_{n}$. We denote by $(\tau,\textup{Disp})$ the $-\mathbb{N}$-algebraic stack induced by the $\tau_{n}$.
Lemma 4.7 ([Bro18, Lem. 2.2.2]).
Let $K^{h,d}_{n,m}$ denote the kernel of the projection $G^{h,d}_{n}\rightarrow G^{h,d}_{m}$ for $m<n$ and $\tilde{K}^{h,d}_{n}$ the kernel of the projection $G^{h,d}_{n}\rightarrow\operatorname{GL}_{h-d}\times\operatorname{GL}_{d}$. Then $K^{h,d}_{n,m}$ and $\tilde{K}^{h,d}_{n}$ are split unipotent.
Remark 4.8.
As seen in the proof of [Bro18, Thm. 2.3.3] the group $G_{1}^{h,d}$ is even a split extension of $\operatorname{GL}_{h-d}\times\operatorname{GL}_{d}$ by a split unipotent group $U_{1}$. The splitting is induced by the canonical inclusion $\operatorname{GL}_{h-d}\times\operatorname{GL}_{d}\hookrightarrow G^{h,d}_{1}$.
Proposition 4.9.
The $-\mathbb{N}$-algebraic stack $(\tau,\textup{Disp})$ is strict monoidal.
Proof.
As DM preserves finite products, it is enough to prove that for any $0\leq d\leq h$ and any $n>m$ the truncation $\tau^{h,d}_{n,m}\colon\textup{Disp}^{h,d}_{n}\rightarrow\textup{Disp}^{h,d}_{m}$ is strict. By definition $\tau_{n,m}$ factors through
[…]
The map $b^{*}\colon\textup{DM}(\left[X^{h}_{m}/G^{h,d}_{m}\right])\rightarrow\textup{DM}(\left[X^{h}_{m}/G^{h,d}_{n}\right])$ is an equivalence of categories by Lemma 4.7 (cf. [RS20, Prop. 2.2.11]).
We now claim that $a^{*}$ is fully faithful (the idea is the same as in [Bro18, Thm 2.3.1]). By definition the map $X^{h}_{n}\rightarrow X^{h}_{m}$ is a $K^{h,0}_{n,m}$-torsor. By étale descent we may replace the morphism $a$ by the projection $K^{h,0}_{n,m}\rightarrow S$. By Lemma 4.7 this is group is split unipotent, so it admits a normal series with successive quotients given by a vector bundle. Thus, applying again étale descent, we may assume that $K^{h,0}_{n,m}$ is a vector bundle over $S$, showing that $a^{*}$ is fully faithful. ∎
Corollary 4.10.
Let $f\colon\textup{Disp}\rightarrow S$ denote the structure map. Then
$f_{\sharp}f^{*}1_{\textup{Disp}}\simeq f_{1\sharp}f_{1}^{*}1_{\textup{Disp}_{1}}\textup{ and }f_{*}f^{*}1_{\textup{Disp}}\simeq f_{1*}f_{1}^{*}1_{\textup{Disp}_{1}}$
are contained in the ind-completion of the full stable subcategory of $\textup{DM}(S)$ generated by Tate-motives.
Proof.
The equivalences follows from Remark 3.2. The fact, that this motive is Artin-Tate follows from Remark 4.8 together with [Yay23a, Lem. 3.12, Thm. 5.1] (note that it is enough to show the assertion for $\textup{Disp}_{1}^{h,d}$). ∎
Remark 4.11.
Brokemper computes the Chow groups of $\textup{Disp}_{1}^{h,d}$ explicitly in [Bro18, Thm. 2.3.3]. Thus, Corollary 4.10 and Remark 3.2 tell us that the motivic cohomology of Disp can be computed by the Chow groups of $\textup{Disp}_{1}^{h,d}$. To be more precise, we can compute
$\bigoplus_{n\in\mathbb{Z}}H^{n,2n}(\textup{Disp},\mathbb{Q})\cong\bigoplus_{0\leq d\leq h}\mathbb{Q}[t_{1},\dots,t_{h}]^{S_{h}\times S_{h-d}}/(c_{1},\dots,c_{h}),$
where $c_{i}$ denotes the $i$-th elementary symmetric polynomial in variables $t_{1},\dots,t_{h}$.
Remark 4.12.
Let BT be the stack of Barsotti-Tate groups (short, BT-groups) over $S$. For any $n\in\mathbb{N}$, we denote by $\textup{BT}_{n}$ the stack of level-$n$ BT-groups. For any $n\geq 0$ there exists a truncation map $\tau_{n}\colon\textup{BT}_{n+1}\rightarrow\textup{BT}_{n}$ and $\lim_{n,\tau}\textup{BT}_{n}\simeq\textup{BT}$. The map $\tau_{n}$ is smooth (cf. [Ill85, Thm. 4.4]) and so the $-\mathbb{N}$-algebraic stack $(\tau,\textup{BT})$ is monoidal. Note that every level-$n$ BT-group admits the notion of a height and dimension, which are locally constant functions over $S$. In particular, we can write $\textup{BT}_{n}\simeq\coprod_{0\leq d\leq h}\textup{BT}^{h,d}_{n}$, where $\textup{BT}^{h,d}_{n}$ denotes the substack of BT generated by BT-groups of dimension $d$ and height $h$.
We want to remark, that there is a morphism $\phi\colon\textup{BT}\rightarrow\textup{Disp}$ compatible with truncations (cf. [Lau13]). Moreover, each of the maps $\phi_{n}$ is smooth and an equivalence on geometric points (cf. [Lau13, Thm. A]). In particular, $\phi_{n}$ is a universal homeomorphism. Note however, that this is not enough to see that $\textup{DM}(\tau,\textup{Disp})\simeq\textup{DM}(\tau,\textup{BT})$ as the morphisms $\phi_{n}$ are not representable. But we think that $\textup{DM}(\tau,\textup{BT})$ is strict monoidal (monoidality follows by a result of Grothendieck [Ill85, Thm. 4.4]). Also we conjecture that $\phi^{*}\colon\textup{DM}(\textup{Disp})\rightarrow\textup{DM}(\textup{BT})$ restricts to Tate-motives proving that the underlying motive of BT and Disp are equivalent.
Remark 4.13.
In another article we will use similar methods to compute the motive of the stack of local $G$-shtukas. Moreover, we will give an explicit description of its absolute motivic cohomology and show that the *-pull/push of the unit is Tate (or to be more precise in the ind-completion of Tate motives).